### v1

# Detecting Discrepancies Between AI-Generated and Natural Images Using Uncertainty

Abstract
In this work, we propose a novel approach for detecting AI-generated images by leveraging predictive uncertainty to mitigate misuse and associated risks. The motivation arises from the fundamental assumption regarding the distributional discrepancy between natural and AI-generated images. The feasibility of distinguishing natural images from AI-generated ones is grounded in the distribution discrepancy between them. Predictive uncertainty offers an effective approach for capturing distribution shifts, thereby providing insights into detecting AI-generated images. Namely, as the distribution shift between training and testing data increases, model performance typically degrades, often accompanied by increased predictive uncertainty. Therefore, we propose to employ predictive uncertainty to reflect the discrepancies between AI-generated and natural images. In this context, the challenge lies in ensuring that the model has been trained over sufficient natural images to avoid the risk of determining the distribution of natural images as that of generated images. We propose to leverage large-scale pre-trained models to calculate the uncertainty as the score for detecting AI-generated images. This leads to a simple yet effective method for detecting AI-generated images using large-scale vision models: images that induce high uncertainty are identified as AI-generated. Comprehensive experiments across multiple benchmarks demonstrate the effectiveness of our method.
1 INTRODUCTION
Recent advancements in generative models have revolutionized image generation, enabling the production of highly realistic images (Midjourney, ; Wukong, ; Rombach et al., 2022). Despite the remarkable capabilities of these models, they pose significant challenges, particularly the rise of deepfakes and manipulated content. The high degree of realism achievable by such technologies prompts urgent discussions about their potential misuse, especially in sensitive domains such as politics and economics. In response to these critical concerns, a variety of methodologies for detecting generated images have emerged. A prevalent strategy treats this detection task as a binary classification problem, necessitating the collection of extensive datasets comprising both natural and AI-generated images to train classifiers (Wang et al., 2020).
While existing detection methods have demonstrated notable successes, they typically encounter challenges in generalizing to images produced by previously unseen generative models Wang et al. (2023a). One promising avenue to enhance the robustness of detection capabilities involves constructing more extensive training datasets by accumulating a diverse array of natural and synthetic images. However, these attempts are often computationally intensive, requiring substantial datasets for effective binary classification. Additionally, maintaining robust detection necessitates continually acquiring images generated by the latest models. And when the latest generative models are not open-sourced, acquiring a large number of generated images to train classifiers is challenging. This highlights the urgent need for a novel framework to detect AI-generated images without requiring AI-generated images.
Figure 1: Large-scale fundamental models trained on a large number of real images are capable of distinguishing between real and generated images.
A recent work (Tan et al., 2024) shows that features extracted by ViT of CLIP (Radford et al., 2021) can be employed to separate natural and AI-generated images, motivating an effective approach to detecting images by training a binary classifier in the feature space of CLIP. This provides a promising direction to explore the possibility that large-scale foundational models already have the ability to capture the subtle differences between real images and AI-generated images. This may be similar to the emergence phenomenon in large language models (Wei et al., 2022), even if these models are not designed to distinguish between real and generated images. This phenomenon may arise from the training data of large vision models. In particular, since training models with AI-generated data usually leads to model collapse (Shumailov et al., 2024), current large vision models are mainly trained over natural data, making these models biased between natural and AI-generated data. Even though natural and AI-generated images appear extremely semantically similar to humans, such “biased” large models can capture the difference between natural and generated images.
The distribution discrepancy in features between natural and generated images motivates us to revisit the strategy of detecting AI-generated images. Specifically, the ability of humans to distinguish between natural and generated images relies on the existence of the discrepancy between them. Moreover, humans are not trained to distinguish natural and generated images. To verify this intuition, we visualize the features of natural and generated images using a large-scale visual fundamental model (DINOv2), following previous work (Tan et al., 2024). As shown in Figure 1, even for images sampled from the same class, there are large distributional discrepancies in the feature space of DINOv2.
Building on this foundation, we propose to leverage predictive uncertainty as the score for distinguishing between natural and AI-generated images. This is because predictive uncertainty offers an effective approach for capturing distribution shifts. In particular, recent studies (Snoek et al., 2019; Schwaiger et al., 2020) indicate that models tend to show increased uncertainty for out-of-distribution (OOD) samples. Hence, for large vision models trained only on natural images, we can treat the natural images as in-distribution samples and the generated images as OOD samples. The challenge comes from efficiently obtaining the uncertainty of the model on the test samples. Classical approaches include Monte-Carlo Dropout (MC-Dropout) (Gal & Ghahramani, 2016) and Deep Ensembles (Lakshminarayanan et al., 2017). However, in our attempts, MC-Dropout obtains sub-optimal results (in Table 5), and it is not practical to train multiple large models independently for ensemble. Instead, we find that this uncertainty can be well captured by perturbing the weights of the models. As shown in Figure 2, when a moderate level of perturbation is applied, the real image has consistent features on the model before and after the perturbation, but the generated image has large differences in features on the model before and after the perturbation.
In this paper, we propose a novel method for AI-generated image detection by weight perturbation (WePe). Our hypothesis is that the model has greater uncertainty in predicting the OOD sample compared to the ID sample, and that this uncertainty can be expressed through sensitivity to weight perturbations. For a large model trained on a large number of real images, the real images can be considered ID samples, while the generated images are considered OOD samples. Thus, the sensitivity of the samples to the weight perturbation of the large model can be an important indicator to determine whether the generative models generate the samples. Despite its simplicity, WePe achieves state-of-the-art performance on various benchmarks.
Figure 2: Natural and generated images are differently sensitive to perturbations in the model weights. When a moderate perturbation is applied ($0.1$), the natural image has essentially the same features on the model before and after the perturbation, but the generated image already shows a large difference.
We summarize our main contributions as follows:
  • We provide a new perspective to detect AI-generated images by calculating predictive uncertainty. This is built upon an intuitive assumption that the natural and generated images differ in distribution, making it possible to employ uncertainty to represent the distribution discrepancy.
  • We propose to leverage large vision models to calculate the predictive uncertainty. The intuition is that large vision models are merely trained on natural images, making it possible to exhibit different uncertainties about natural and generated images. We capture this uncertainty by weight perturbation to effectively detect images (Eq. 3).
  • Comprehensive experiments on multiple benchmarks demonstrate that the proposed method outperforms previous methods, including training-based methods, achieving state-of-the-art performance.
2 Related Works
AI-Generated images detection. Recent advancements in generative models, such as those by (Brock et al., 2019; Ho et al., 2020), have led to the creation of highly realistic images, highlighting the urgent need for effective algorithms to distinguish between natural and generated images. Prior research, including works by (Frank et al., 2020; Marra et al., 2018), primarily focuses on developing specialized binary classification neural networks to differentiate between natural and generated images. Notably, CNNspot (Wang et al., 2020) demonstrates that a standard image classifier trained on ProGAN can generalize across various architectures when combined with specific data augmentation techniques. F3Net (Qian et al., 2020) distinguishes real face images from fake face images with the help of frequency statistical differences. NPR (Tan et al., 2024) introduces the concept of neighboring pixel relationships to capture differences between natural and generated images. Although these methods show superior performance on generators in the training set, they often do not generalize well to unknown generators. In addition to this, training-based methods are susceptible to small perturbations in the image. For this reason, recently, some training-free methods have been proposed. AEROBLADE (Ricker et al., 2024) calculates the reconstruction error with the help of the autoencoder used in latent diffusion models (Rombach et al., 2022). RIGID (He et al., 2024) finds that real images are more robust to small noise perturbations than generated images in the representation space of the vision foundation models and exploits this property for detection. However, these methods usually make overly strong assumptions about natural or generated images, leading to insufficient generalization. In our paper, we propose a training-free detection method through uncertainty analysis. Based on the widespread phenomenon that generated images have greater uncertainty than real images on models trained with real images, our method achieves robust detection performance.
Uncertainty estimation. Uncertainty estimation in machine learning has seen significant advancements in recent years. (Gal & Ghahramani, 2016) introduces Monte Carlo Dropout (MC Dropout), which uses dropout at inference to estimate uncertainty from the variance of multiple predictions. (Lakshminarayanan et al., 2017) develops deep ensembles, demonstrating improved uncertainty estimates through training multiple model indepently with different initializations. Recent work by (Snoek et al., 2019) analyzes the calibration of uncertainty in deep learning models, highlighting the importance of reliable uncertainty measures. Additionally, (Guo et al., 2017) explore the use of temperature scaling to enhance the calibration of model predictions. In our paper, We measure uncertainty by perturbing the model’s weights.
Random weight perturbations. Weight perturbation, i.e. adding noise to network weights, has been extensively studied. Many methods (Khan et al., 2018; Wu et al., 2020) perform training regularization by adding noise to the weights of the neural network during training to improve the generalization of the network. Weight perturbation has also been applied in adversarial attacks to study the robustness of the network (He et al., 2019; Garg et al., 2020). In addition to this, some works also study the sensitivity of neural network to weights perturbation. (Cheney et al., 2017) shows that convolutional networks are surprisingly robust to a number of internal perturbations in the higher convolutional layers but the bottom convolutional layers are much more fragile. (Weng et al., 2020) proposes an efficient approach to compute a certified robustness bound of weight perturbations. In our work, we find neural networks trained on natural images exhibit different robustness to weight perturbations for natural and generated images. Based on this property, we propose an efficient algorithm to distinguish between natural and generated images.
3 Method
3.1 Motivation
Our method is built upon a foundational assumption that natural and generated images have different distributions. This is a reasonable assumption; otherwise, we cannot distinguish between natural and generated images. Fortunately, this assumption is consistent with previous work (Tan et al., 2024) and our empirical observations, as shown in Figure 1.
Hence, for the large models trained merely on natural images, we can regard natural images as in-distribution (ID) data while generated images as out-of-distribution (OOD) data. This distribution discrepancy can be reflected by the widely used predictive uncertainty, since neural networks typically exhibit higher uncertainty for OOD samples (Snoek et al., 2019; Schwaiger et al., 2020). This leads to a simple yet novel approach to determine whether a test image is generated by AI models when we can calculate its uncertainty on a pre-trained large vision model. In the following, we first give a complete introduction to our method WePe, then describe on our reasons for choosing the method, and finally, we explore why WePe works.
3.2 Uncertainty Definition
Classical methods of uncertainty estimation, such as Deep Ensembles and MC Dropout, can simply be viewed as using the variance of the results of multiple predictions as an estimation of uncertainty $u(\mathbf{x})$:
$\mu(\mathbf{x})=\frac{1}{n}\sum_{t=1}^{n}\hat{y}_{t}(\mathbf{x}),\quad u(\mathbf{x})=\sigma^{2}=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{y}_{t}(\mathbf{x})-\mu(\mathbf{x})\right)^{2},$ (1)
where, $\hat{y}_{t}$ denotes the $t$-th prediction. The multiple predictions of Deep Ensembles come from multiple independently trained neural networks, while the multiple predictions of MC Dropout come from the use of dropout during inference, which can be regarded as multiple prediction using neural networks with different structures.
3.3 Uncertainty Calculation
The predictive uncertainty is typically calculated as the variance of predictions obtained with certain perturbations. In this work, we simply leverage $\theta$ as the features or parameters before perturbation. Specifically, the predictive uncertainty $u(\mathbf{x})$ can be calculated by,
$u(\mathbf{x})=\frac{1}{n}\sum_{k=1}^{n}[f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta_{t})-\sum_{j=1}^{n}\frac{f(\mathbf{x};\theta_{j})^{\top}f(\mathbf{x};\theta_{t})}{n}]^{2},$ (2)
where $n$ is the number of samples, $f(\mathbf{x};\theta_{k})$ denotes the L2-normalized features of an input image $\mathbf{x}$ when inferring with the parameter $\theta_{k}$, and $\theta_{t}$ stands for the teacher model used in DINOv2.
However, we cannot access the teacher model $\theta_{t}$, making it challenging to calculate the uncertainty. Moreover, even if it is available, introducing two models for calculation leads to low computation efficiency. Fortunately, we can calculate an upper bound of $u(\mathbf{x})$. This can be formalized by,
$u(\mathbf{x})\leq\frac{1}{n}\sum_{k=1}^{n}\left\|f(\mathbf{x};\theta_{k})-\frac{1}{n}\sum_{j}^{n}f(\mathbf{x};\theta_{j})\right\|^{2}\left\|f(\mathbf{x};\theta_{t})\right\|^{2}=2-\frac{2}{n}\sum_{k=1}^{n}f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta),$ (3)
where $\theta$ denotes the parameter before injecting perturbation, and we leverage an unbiased assumption that the expectation $\mathbb{E}_{\theta_{j}}f(\mathbf{x};\theta_{j})$ approaches the feature $f(\mathbf{x};\theta)$ extracted by the non-perturbed parameter $\theta$. Eq. 3 provides a simple approach to calculate the uncertainty without needing a teacher model used in the training phase of DINOv2. The insight of Eq. 3 is intuitive. Specifically, if an image $\mathbf{x}$ causes a high feature similarity between the original and perturbed parameter, the image leads to a low uncertainty and is more likely to be a natural image.
3.4 An overview of WePe
As discussed above, the proposed WePe is based on a large model pre-trained on a large number of natural images. In this work, we chose DINOv2 (Oquab et al., 2024), a large model trained with contrastive learning on image data. In order to capture the uncertainty of the model on the test images, we extract image features using the original model and the model after adding noise to the parameters respectively. The similarity between the pre-perturbation and post-perturbation feature vectors is quantified using a suitable distance metric, such as cosine similarity. Images exhibiting high similarity are classified as real, while those with low similarity are identified as generated. This method not only capitalizes on the characteristics of the DINOv2 model but also provides a robust framework for distinguishing between real and generated images based on their feature stability under model perturbation.
3.5 Discussions
Why choose DINOv2? In addition to DINOv2, CLIP is a commonly used model. However, our experiments show that CLIP performs sub-optimally compared to DINOv2 (see Table 5). We believe this difference stems from their training strategies. Unlike DINOv2, CLIP is a multimodal model that combines image and textual features from captions for contrastive learning, which may lead it to focus on broader semantic features rather than fine details. In contrast, DINOv2 emphasizes contrastive learning solely on images, allowing it to better capture subtle differences between natural and generated images. Therefore, we use DINOv2 in our experiments.
Why choose weight perturbation? Common methods for measuring uncertainty include MC Dropout and Deep Ensembles. MC Dropout involves keeping dropout active during testing and performing multiple forward passes on the inputs to generate outputs with different network structures. The variability among these outputs serves as an estimate of the model’s uncertainty regarding the input data. However, since DINOv2 does not utilize dropout during training, MC Dropout may not yield optimal results (see Table 5). Deep Ensembles, on the other hand, trains multiple networks independently and uses the differences in their outputs on test samples to assess uncertainty. However, training multiple DINOv2-level models is impractical. Therefore, in our study, we choose to perturb the model parameters and assess the differences in outputs from the original and perturbed models to estimate uncertainty for the test images.
Why does weight perturbation work? Incorporating weight perturbations during testing can effectively simulate Bayesian inference, thereby capturing the inherent uncertainty in neural networks. From a Bayesian perspective, the weights of a neural network are not fixed but rather distributions reflecting the range of plausible values given the data. By introducing noise into the network weights at test time, we mimic the process of drawing samples from a posterior distribution over weights, a core concept in Bayesian inference. This technique enables the model to generate diverse predictions, reflecting its uncertainty, particularly when encountering out-of-distribution samples. This approach is akin to Bayesian neural networks, where weight uncertainty is explicitly modeled. For example, (Blundell et al., 2015) proposed using variational methods to approximate weight distributions in Bayesian neural networks, allowing for uncertainty quantification through weight perturbations. Similarly, (Gal & Ghahramani, 2016) demonstrated that introducing dropout at test time serves as a Bayesian approximation, with the added noise acting as a proxy for weight sampling, thus allowing for reliable uncertainty estimation. In this paper, we capture uncertainty by adding noise directly to the model parameters to distinguish between natural and generated images. However, trying to theoretically analyze the different sensitivities of natural and generated images to model weights in the representation space of a large model is a difficult task. Therefore, we only draw this conclusion through various empirical observations in this paper, and do not directly prove this theoretically. The theoretical proof will be left to our future work.
4 Experiments
4.1 Experiment setup
Datasets. Following previous work (He et al., 2024), we evaluate the performance of WePe on ImageNet (Deng et al., 2009), LSUN-BEDROOM (Yu et al., 2015) and GenImage (Zhu et al., 2023). For ImageNet and LSUN-BEDROOM, the generated images are provided by  (Stein et al., 2023). For ImageNet, the generative models include ADM (Dhariwal & Nichol, 2021), ADM-G, LDM (Rombach et al., 2022), DiT-XL2 (Peebles & Xie, 2023), BigGAN (Brock et al., 2019), GigaGAN (Kang et al., 2023), StyleGAN (Karras et al., 2019), RQ-Transformer (Lee et al., 2022), and MaskGIT (Chang et al., 2022). For LSUN-BEDROOM,the generative models include ADM, DDPM (Ho et al., 2020), iDDPM (Nichol & Dhariwal, 2021), Diffusion Projected GAN (Wang et al., 2023b), Projected GAN (Wang et al., 2023b), StyleGAN (Karras et al., 2019) and Unleasing Transformer (Bond-Taylor et al., 2022). GenImage primarily employs the Diffusion model for image generation. The fake images are generated by Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), GLIDE, VQDM (Gu et al., 2022), Wukong (Wukong, ), BigGAN, ADM, and Midjourney (Midjourney, ). It contains 1,331,167 real and 1,350,000 generated images.
Evaluation metrics. Following RIGID, we mainly use the following metrics: (1) the average precision (AP), (2) the area under the receiver operating characteristic curve (AUROC). On GenImage, since the dataset is too large and re-implementing all baselines is time-consuming, we obtain the results directly from the corresponding papers and report the classification accuracy (ACC).
Baselines. Following RIGID, we take both training-free methods and training methods as baselines. For training-free methods, we take RIGID (He et al., 2024) and AEROBLADE (Ricker et al., 2024) as baselines. For training methods, we take DIRE (Wang et al., 2023a), CNNspot (Wang et al., 2020), Ojha (Ojha et al., 2023) and NPR (Tan et al., 2024) as baselines. Besides, on GenImage, we also report the result of Frank (Frank et al., 2020), Durall (Durall et al., 2020), Patchfor (Chai et al., 2020), F3Net (Qian et al., 2020), SelfBland (Shiohara & Yamasaki, 2022), GANDetection (Mandelli et al., 2022), LGrad (Tan et al., 2023), ResNet-50 (He et al., 2016), DeiT-S (Touvron et al., 2021), Swin-T (Liu et al., 2021), Spec (Zhang et al., 2019), GramNet (Liu et al., 2020). For these baselines, we get the results directly in the corresponding papers without reimplementing them.
Experiment details. To balance detection performance and efficiency, we use DINOv2 ViT-L/14. We report the average results under five different random seeds and report the variance in Figure 8. In our experiments we find that perturbing the high layers may lead to a large corruption in the features of the real images, resulting in sub-optimal results. Therefore, We do not perturb the high-level parameters. In DINOv2 ViT-L/14, the model has 24 transformer blocks, and we only perturb the parameters of the first 19 blocks with Gaussian perturbations of zero mean. The variance of the added Gaussian noise is proportional to the mean value of the parameters in each block, with the ratio set to 0.1. Ablation experiments show that our method is robust to the blocks chosen for perturbation and the level of noise.
4.2 Results
Comparison with other baselines. We conduct full comparative experiments on the three benchmarks mentioned. As shown in Table 1,  2 and  3, WePe achieves the best detection performance on ImageNet, LSUN-BEDROOM and GenImage without the need for training. It is worth noting that on the large-scale GenImage benchmark, the training-based method, despite having perfect performance on the generators used during training, performs poorly on many generators not seen during training. This illustrates the extreme dependence of the performance of training-based methods on the diversity of the training set. In contrast, our method does not require training, performs well on a wide variety of generators, and outperforms the SOTA training method by $3.36\%$ on average. On generators that have not been seen during training, such as VQDM, many training methods exhibit random prediction results, whereas WePe achieves superior detection performance. To further illustrate the effectiveness of our method, we count the difference in feature similarity between real and fake images on the pre- and post-perturbation models. As shown in Figure 3, the small perturbation of the model has less effect on the real images than on generated images, resulting in higher feature similarity before and after the perturbation. The discrepancy effectively distinguishes the real image from the generated image.
Figure 3: Comparison of cosine similarity between original and transformed images. The generated images are from: (a) ADM, (b) BigGAN, and (c) DDPM.
Table 1: AI-generated image detection performance on ImageNet. Values are percentages. Bold numbers are superior results and the underlined italicized values are the second-best performance. A higher value indicates better performance.
Models
ADM ADMG LDM DiT BigGAN GigaGAN StyleGAN XL RQ-Transformer Mask GIT Average
Methods AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP
Training Methods
CNNspot 62.25 63.13 63.28 62.27 63.16 64.81 62.85 61.16 85.71 84.93 74.85 71.45 68.41 68.67 61.83 62.91 60.98 61.69 67.04 66.78
Ojha 83.37 82.95 79.60 78.15 80.35 79.71 82.93 81.72 93.07 92.77 87.45 84.88 85.36 83.15 85.19 84.22 90.82 90.71 85.35 84.25
DIRE 51.82 50.29 53.14 52.96 52.83 51.84 54.67 55.10 51.62 50.83 50.70 50.27 50.95 51.36 55.95 54.83 52.58 52.10 52.70 52.18
NPR 85.68 80.86 84.34 79.79 91.98 86.96 86.15 81.26 89.73 84.46 82.21 78.20 84.13 78.73 80.21 73.21 89.61 84.15 86.00 80.84
Training-free Methods
AEROBLADA 55.61 54.26 61.57 56.58 62.67 60.93 85.88 87.71 44.36 45.66 47.39 48.14 47.28 48.54 67.05 67.69 48.05 48.75 57.87 57.85
RIGID 87.16 85.08 80.09 77.07 72.43 69.30 70.40 65.94 90.08 89.26 86.39 84.11 86.32 85.44 90.06 88.74 89.30 89.25 83.58 81.58
WePe 89.79 87.32 83.20 78.80 78.47 73.50 77.13 71.21 94.24 93.64 92.15 90.29 93.86 92.86 93.50 91.47 89.55 86.25 87.99 85.04
Table 2: AI-generated image detection performance on LSUN-BEDROOM.
Models
ADM DDPM iDDPM Diffusion GAN Projected GAN StyleGAN Unleashing Transformer Average
Methods AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP
CNNspot 64.83 64.24 79.04 80.58 76.95 76.28 88.45 87.19 90.80 89.94 95.17 94.94 93.42 93.11 84.09 83.75
Ojha 71.26 70.95 79.26 78.27 74.80 73.46 84.56 82.91 82.00 78.42 81.22 78.08 83.58 83.48 79.53 77.94
DIRE 57.19 56.85 61.91 61.35 59.82 58.29 53.18 53.48 55.35 54.93 57.66 56.90 67.92 68.33 59.00 58.59
NPR 75.43 72.60 91.42 90.89 89.49 88.25 76.17 74.19 75.07 74.59 68.82 63.53 84.39 83.67 80.11 78.25
AEROBLADA 57.05 58.37 61.57 61.49 59.82 61.06 47.12 48.25 45.98 46.15 45.63 47.06 59.71 57.34 53.85 54.25
RIGID 71.90 72.29 88.31 88.55 84.02 84.80 91.42 91.90 92.12 92.54 77.29 74.96 91.37 91.39 85.20 85.20
WePe 73.85 70.21 88.84 87.14 86.23 83.82 94.16 93.52 95.34 95.18 83.50 80.66 94.18 93.45 88.01 86.28
Table 3: AI-generated image detection performance on GenImage. Except for WePe and RIGID, all methods require training on SD V1.4. The results of the baseline method are all from GenImage
Models
Methods Midjourney SD V1.5 ADM GLIDE Wukong VQDM BigGAN Average
ResNet-50 54.90 99.70 53.50 61.90 98.20 56.60 52.00 68.11
DeiT-S 55.60 99.80 49.80 58.10 98.90 56.90 53.50 67.51
Swin-T 62.10 99.80 49.80 67.60 99.10 62.30 57.60 71.19
CNNspot 52.80 95.90 50.10 39.80 78.60 53.40 46.80 58.63
Spec 52.00 99.20 49.70 49.80 94.80 55.60 49.80 64.41
F3Net 50.10 99.90 49.90 50.00 99.90 49.90 49.90 64.22
GramNet 54.20 99.10 50.30 54.60 98.90 50.80 51.70 65.66
DIRE 60.20 99.80 50.90 55.00 99.20 50.10 50.20 66.49
Ojha 73.20 84.00 55.20 76.90 75.60 56.90 80.30 71.73
LaRE 66.40 87.10 66.70 81.30 85.50 84.40 74.00 77.91
RIGID 81.54 68.72 72.35 84.15 68.57 78.98 93.02 78.19
WePe 79.17 75.57 76.07 79.20 79.00 90.60 89.27 81.27
Figure 4: Performance varies with variance
Figure 5: The effects of disturbed blocks.
Figure 6: Illustration of detection performance varying with perturbation intensity under different degradation mechanisms, (a): JPEG compression, (b): Gaussian blur, and (c): Gaussian noise.
4.3 Ablation study
In this section, we perform ablation experiments. Unless otherwise stated, experiments are conducted on ImageNet benchmark.
Robustness to Perturbations. Robustness to various perturbations is a critical metric for detecting generated images. In real-world scenarios, images frequently undergo perturbations that can impact detection performance. Following RIGID, we assess the robustness of detectors against three types of perturbations: JPEG compression (with quality parameter $q$)), Gaussian blur (with standard deviation $\sigma$), and Gaussian noise (with standard deviation $\sigma$). As illustrated in Figure 6, training-free methods generally exhibit superior robustness compared to training-based methods, with our method achieving the best overall performance.
The impact of the degree of perturbation. As shown in Figure 5, we explore the effect of the degree of perturbation on the performance of WePe. It can be seen that WePe is quite robust to the level of perturbation noise. It is only when the noise is very large or very small that it leads to a degradation in performance. When the noise level is small, the features obtained before and after the model perturbation are extremely similar, while when the noise level is very large, the features obtained before and after the model perturbation are extremely dissimilar, and these two cases will result in the inability to effectively differentiate between real and generated images.
Selecting which layers’ parameters to perturb? As shown in Figure 5, we explore the choice of which layers’ parameters to perturb would achieve good performance. The horizontal coordinates in the graph indicate that the first $k$ blocks are perturbed, not the $kth$ block. The experimental results exhibit that our method obtains good performance when the parameters of the first 9 to the first 20 blocks are chosen to be perturbed. This demonstrates the robustness of our method. In practice, we can select the layers to be perturbed by a small set of real and generated images. And when the generated images are not available, we can also use the probe to determine which layers are perturbed using only the real image. We describe our method in Appendix A.14.
The effect of models. In our experiments, we mainly used DINOv2 ViT-L/14 to extract features. We further explore the effect of using other models of DINOv2, including ViT-S/14, ViT-B/14, and ViT-g/14. In addition to this, we conduct experiments on the CLIP:ViT-L/14. As shown in Table 5, the performance on CLIP is not as good as on DINOv2. We hypothesize that the difference comes from the training approach of these models. CLIP learns features using image captions as supervision, which may make the features more focused on semantic information, whereas DINOv2 learns features only from images, which makes it more focused on the images themselves, and thus better able to capture subtle differences in real and generated images. When using DINOv2, our methods performs poorly when the capacity of the model is small. This may be due to the fact that larger models can better capture the differences between real and fake images. In our main experiments, to balance detection efficiency and detection performance, we use DINOv2 ViT-L/14.
The effect of perturbation type. In our experiments, we perturb the model parameters by adding Gaussian noise. We further explore other ways of perturbation, such as adding uniform noise or Laplace noise to the weight. In addition to this, we also explore the effect of MC Dropout, i.e., using dropout during inference. As shown in Table 5, all three weight perturbation methods achieve good performance, and outperform MC Dropout.
Table 4: The effect of models.
model      AUROC      AP DINOv2: ViT-S/14      72.83      71.63 DINOv2: ViT-B/14      81.82      80.64 DINOv2: ViT-L/14      87.99      85.04 DINOv2: ViT-g/14      84.92      81.83 CLIP: ViT-L/14      77.89      77.90
Table 5: The effect of type of perturbation.
noise      AUROC      AP Gaussian noise      87.99      85.04 Uniform noise      89.06      86.32 Laplace noise      87.13      84.22 MC Dropout      81.63      79.71
5 Limitation
The proposed weight perturbation provides a simple and effective method for detecting generated images, yet we have not theoretically justified the widespread use of the method due to the inclusion of a variety of strong prior assumptions, such as the assumption about treating generated samples as OOD data. Therefore, our future work will focus on establishing the theoretical foundations of our method.
6 Conclusion
In this work, to effectively address the challenges of detecting AI-generated images, we propose a novel approach that leverages predictive uncertainty as a key metric. Our findings reveal that by analyzing the discrepancies in distribution between natural and AI-generated images, we can significantly enhance detection performance. The use of large-scale pre-trained models allows for accurate computation of predictive uncertainty, enabling us to identify images with high uncertainty as likely AI-generated. Our method achieves robust detection performance in a simple untrained manner. Overall, our approach demonstrates a promising direction for improving AI-generated image detection and mitigating potential risks associated with their misuse. Future work could delve deeper into refining the predictive models and exploring additional features that could further enhance detection accuracy.
ETHIC STATEMENT
According to the Code of Ethics, our work does not raise any ethical concerns, as there are no human subjects, private datasets, harmful insights, or research integrity issues. Our work does not raise any ethical concerns. In particular, our work focuses on detecting AI-generated images to avoid the potential risk induced by the development of generative models.
REPRODUCIBILITY STATEMENT
To highlight the efforts that have been made to ensure reproducibility, we summarize materials facilitating reproducible results:
  • Theoretical results. Our work mainly focuses on the empirical investigation without explicit theatrical results.
  • Datasets. All involved datasets are publicly available. To facilitate reproducible results, we provide details in Sec 4 and Appendix A.15.
  • Open Source. We will release our code once the paper is accepted. If reviewers request, we will include an anonymized link to the code in our response.
Appendix A Appendix
A.1 Concerns about models being contaminated by generated images
With the proliferation of generated images, future large-scale models could become contaminated by such data, making it increasingly difficult to distinguish between natural and generated images. One potential solution is to employ machine unlearning (Yao et al., 2023) techniques to detect and address generated images. Machine unlearning focuses on removing the influence of specific data from pre-trained models, primarily due to privacy concerns. In this context, when generative images are incorporated into the training process of large-scale models, we can utilize machine unlearning techniques to eliminate the effects of these images on pre-trained models. This approach would help ensure that the features of natural and generated images remain distinct and separable.
A.2 Discussion on distribution discrepancy
In this paper, the core assumption we make is that there is data distribution discrepancy between natural and generated images. This assumption is valid for current generative models and has been confirmed by many works (Corvi et al., 2023; Tan et al., 2024; Ricker et al., 2024). This assumption is also the foundation of many generative image detection methods (we cannot distinguish between images that are indistinguishable).
Secondly, we observe that this discrepancy in data distribution can be captured by the representation space of a vision model pre-trained on a large number of natural images, i.e., there is feature distribution discrepancy between the generated and natural images, as shown in Figure 7. However, this remains an observation, and we have not found theoretical proof despite reviewing the literature. We only observe a similar phenomenon in UnivFD (Ojha et al., 2023), where the feature distribution discrepancy is observed in the representation space of CLIP:ViT-L/14.
That said, we can confirm the existence of feature distribution discrepancy of generated and natural images based on an important metric for evaluating generative models, the FID score. The FID score measures the feature distribution discrepancy between natural and generated images on the Inception network (Szegedy et al., 2015). When the FID score is 0, it indicates that the two distributions do not differ. However, even on these simple networks such as Inception v3, advanced generative models like ADM still achieve an FID score of 11.84, not to mention that on powerful models such as DINOv2, we observe significant feature distribution discrepancy.
Figure 7: Feature distribution discrepancy between the generated and natural images on DINOv2 and CLIP. $\bullet$ and $\bullet$ represent the feature of natural images and AI-generated images on the corresponding models.
A.3 Measuring feature distribution discrepancy with FID scores
We further use the ”FID” score to measure the difference in feature distribution between natural and generated images. To avoid the effects of categories, we compute the FID scores using the DINOv2 model on the LSUN-BEDROOM benchmark. For each category of images, we randomly select 5000 images for calculation. In addition to calculating the FID scores between natural images and generated images, we also calculate the FID scores between natural images and natural images. As shown in Table 6, the FID scores between natural images and generated images are significantly higher than the FID scores between natural images and natural images. Moreover, there is a clear positive correlation between the detection performance of WePe and the FID score. This result fully explains the existence of feature distribution discrepancy between natural and generated images on DINOv2, and demonstrates that WePe can effectively detect the feature distribution discrepancy.
Table 6: Measuring feature distribution discrepancy with FID scores.
Models Natural ADM StyleGAN iDDPM DDPM Diffusion GAN Unleashing Transformer Projected GAN
FID score 1.09 18.25 52.31 59.94 80.44 116.06 130.00 132.75
AUROC 50.00 73.85 83.50 86.23 88.84 94.16 94.18 95.34
A.4 Comparison with Gradient Cuff
Gradient Cuff (Hu et al., 2024) focuses on detecting jailbreak attacks in Large Language Models (LLMs). It finds that the landscape of the refusal loss is more precipitous for malicious queries than for benign queries. And then it uses stochastic gradient estimation to estimate gradient and use the gradient norm as the decision score. Thus, we can also leverage this interesting work to identify the distribution discrepancy for detection. We use the similar way to estimate gradient as the decision score. As shown in Table 8, we surprisingly find that this method even surpasses WePe. This suggests that it is also possible to distinguish between natural and generated images by estimating the gradient.
A.5 WePe on large multi-modal models
In addition to CLIP, we further test the performance of WePe on BLIP (Li et al., 2022). As shown in Table 8, the performance of WePe is unsatisfactory on these multimodal models, which may be due to the fact that the image features of the multimodal models are more focused on semantic information, in line with our discussions.
A.6 Performance on adversarial examples
We further test WePe on adversarial examples. We simply add Gaussian noise (with different standard deviation $\sigma$) to the test samples to simulate the adversarial examples. We test three cases: adding noise on the natural image, adding noise on the generated image and adding noise on all images. As shown in Table 10, when noise is injected, the feature similarity between the clean model and the noisy model for the noisy image decreases, which leads to a change in the detection performance. To mitigate this effect, we can perform detection by perturbing the model multiple times and using the average similarity. The effect of noise is successfully mitigated by ensemble as shown in Table 10.
Table 7: Comparison with Gradient Cuff.
Method      AUROC      AP WePe      87.99      85.04 WePe + Gradient Cuff      89.36      90.62
Table 8: WePe on large multi-modal models.
Model      AUROC      AP DINOv2      87.99      85.04 CLIP      77.89      77.90 BLIP      68.25      64.68
Table 9: Performance on adversarial examples.
Case      AUROC      AP Clean images      87.99      85.04 Noisy natural images, $\sigma=0.05$      81.79      78.99 Noisy natural images, $\sigma=0.1$      76.16      73.16 Noisy natural images, $\sigma=0.15$      69.53      66.83 Noisy generated images, $\sigma=0.05$      88.03      85.82 Noisy generated images, $\sigma=0.1$      88.77      87.14 Noisy generated images, $\sigma=0.15$      90.25      88.99 Noisy natural and generated images, $\sigma=0.05$      84.94      81.98 Noisy natural and generated images, $\sigma=0.1$      81.46      78.25 Noisy natural and generated images, $\sigma=0.15$      78.19      75.32
Table 10: Mitigating the effects of image noise through ensemble.
Case      AUROC      AP Clean images      87.99      85.04 Noisy natural images, $\sigma=0.05,n=1$      81.79      78.99 Noisy natural images, $\sigma=0.05,n=5$      85.30      82.59 Noisy natural images, $\sigma=0.1,n=1$      76.16      73.16 Noisy natural images, $\sigma=0.1,n=5$      81.36      78.39 Noisy natural images, $\sigma=0.15,n=1$      69.53      66.83 Noisy natural images, $\sigma=0.15,n=5$      75.16      72.58
A.7 Concerns about hard samples
WePe relies on the model being pre-trained on a large dataset of natural images. Given the abundance of natural images, it is possible WePe may misclassify other natural images that are out-of-distribution, as AI-generated images, leading to false-negative errors. Thus, we follow previous work to leverage AUROC as the main metric the evaluate different methods. This is because AUROC is a pivotal metric for reflecting the false-positive and false-negative costs. The hard (natural) samples play a crucial role in detecting generated images. We will explore how to leverage hard samples to promote the detection performance in our future work.
A.8 Boosting performance on CLIP
As mentioned above, WePe does not perform satisfactorily on vision language models such as CLIP. Since CLIP needs to unite image features and text features, the image features need to be projected into the union space, which results in the projected image features being more focused on semantic information. For this reason, we can improve the performance of WePe on CLIP by using the features before projection for detection, as shown in Table 12.
Table 11: Boosting performance on CLIP.
Model      AUROC      AP DINOv2      87.99      85.04 CLIP      77.89      77.90 CLIP without projection matrix      84.82      84.20
Table 12: Comparison of detection times.
Method      Time (s) AEROBLADE      17.6 RIGID      3.7 WePe      4.5
A.9 Comparison of detection times.
Table 13: Evading detection of RIGID by adding noise to the generated image.
Model      AUROC      AP
RIGID      83.58      81.58
WePe      87.99      85.04
RIGID with noisy generated images      18.69      34.51
WePe with noisy generated images      88.77      87.14
Our method use a perturbed pre-trained model that is fixed during inferring all test samples. Thus, our method can be processed within two forward passes. This is equal to the cost of RIGID that requires two forward passes for clean and noisy images. However, RIGID can concatenate clean and noisy images in a mini batch and obtain detection results by with a single forward pass. AEROBLADE requires only one forward pass, but it needs to compute the reconstruction error of the image. This takes a long time to reconstruct at the pixel level. Besides, AEROBLADE needs to use a neural network to compute the LPIPS score, leading to computational complexity. As shown in Table 12, we compare the time required to detect 100 images under the same conditions. Since AEROBLADE needs to calculate the image reconstruction error, it has the lowest detection efficiency. RIGID can obtain detection results in a single forward pass by concatenating clean and noisy images, whereas WePe requires two forward passes, which results in WePe’s detection efficiency being inferior to RIGID’s. However, WePe can be parallelized across two devices to obtain the detection results in a single forward pass.
A.10 Comparison with RIGID
The main differences between WePe and RIGID are as follows:
  • The approach proposed by RIGID stems only from the phenomenon it observed: namely, that natural and generated images show different sensitivities to noise in the representation space of DINOv2. Instead, WePe explicitly proposes that there is distribution discrepancy between natural and generated images, and utilizes the difference in uncertainty to expose distribution discrepancy between natural and generated images.
  • RIGID utilizes the difference in sensitivity to noise between natural image and generated image for detection. Although the generated image is more sensitive, it is easy to think of a way to avoid detection, i.e., the user adds noise to the generated image and then submits it for detection. This approach easily removes the sensitivity of the generated image to noise. As shown in Table 13, after adding noise to a generated image, RIGID determines the noisy generated image as natural images. This is a fatal flaw of RIGID. However, WePe is not affected by this. WePe exposes the difference in distribution between the test image and the natural image through weight perturbation. Adding noise to the generated image further increases this difference, leading to even better detection performance.
A.11 Note on the unbiased assumption
In section 3.3, we make an assumption that the expected extracted feature by noised models is unbiased for that by the original model. We think this assumption is reasonable. Thanks to the over-parameterization of modern neural networks and advanced optimization algorithms (e.g., AdamW), it is a well-established fact that trained neural networks are usually smooth in the parameter space and show robustness to small weight perturbation (Novak et al., 2018). And this robustness is used in many applications, such as quantization (Gholami et al., 2022) and pruning (Liu et al., 2019). And in Figure 2, we also clearly show this robustness: when adding tiny noise to the model weights, the features remain almost unchanged.
A.12 SOFTWARE AND HARDWARE
We use python 3.8.16 and Pytorch 1.12.1, and seveal NVIDIA GeForce RTX-3090 GPU and NVIDIA GeForce RTX-4090 GPU.
Figure 8: WePe with multiple perturbations.
A.13 WePe with multiple perturbation
In our experiments, taking into account the detection efficiency, we perturb the model only once, and then calculate the feature similarity of the test samples on the clean and perturbed models. We further experiment with multiple perturbations and use the mean of the feature similarity of the test samples on the clean model and all the perturbed models as the criterion for determining whether or not the image is generated by the generative models. As shown in Figure 8,
A.14 Using natural images only to select which layers to perturb
In our experiments, we use a small set of natural images and generated images to pick the parameters that need to be perturbed. When all the generated images are not available, we can also use only the real images to select the layers that need to be perturbed. Specifically, we first perturb each block alone and calculate the similarity of the features on the model of the natural image before and after the perturbation, as shown in Table 14. We then sort the similarity and select the blocks with the highest similarity for perturbation. As shown in Table 15, selecting the parameters to be perturbed in this way also achieves good performance and has strong robustness.
Table 14: Effect of perturbation position on natural images. We perturb each block individually, observe the similarity of features on the model of the natural image before and after the perturbation and rank these blocks.
block 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23
similarity(%) 99.40 97.66 98.83 99.00 98.80 98.70 99.37 94.73 92.87 98.44 97.07 98.00 93.46 96.24 94.80 93.85 92.40 87.60 71.50 76.00 80.27 75.93 34.81 47.90
rank 1 9 4 3 5 6 2 13 16 7 10 8 15 11 12 14 17 18 22 20 19 21 24 23
Table 15: AI-generated image detection performance on ImageNet. We select the top-k blocks with the highest similarity for perturbation based on the sorting results.
Models
ADM ADMG LDM DiT BigGAN GigaGAN StyleGAN XL RQ-Transformer Mask GIT Average
Methods AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP
Training Methods
CNNspot 62.25 63.13 63.28 62.27 63.16 64.81 62.85 61.16 85.71 84.93 74.85 71.45 68.41 68.67 61.83 62.91 60.98 61.69 67.04 66.78
Ojha 83.37 82.95 79.60 78.15 80.35 79.71 82.93 81.72 93.07 92.77 87.45 84.88 85.36 83.15 85.19 84.22 90.82 90.71 85.35 84.25
DIRE 51.82 50.29 53.14 52.96 52.83 51.84 54.67 55.10 51.62 50.83 50.70 50.27 50.95 51.36 55.95 54.83 52.58 52.10 52.70 52.18
NPR 85.68 80.86 84.34 79.79 91.98 86.96 86.15 81.26 89.73 84.46 82.21 78.20 84.13 78.73 80.21 73.21 89.61 84.15 86.00 80.84
Training-free Methods
AEROBLADA 55.61 54.26 61.57 56.58 62.67 60.93 85.88 87.71 44.36 45.66 47.39 48.14 47.28 48.54 67.05 67.69 48.05 48.75 57.87 57.85
RIGID 87.16 85.08 80.09 77.07 72.43 69.30 70.40 65.94 90.08 89.26 86.39 84.11 86.32 85.44 90.06 88.74 89.30 89.25 83.58 81.58
WePe top-8 89.25 86.53 82.66 78.08 79.29 73.88 78.53 72.48 93.90 92.61 92.07 89.65 93.06 91.26 92.68 89.84 89.85 86.91 87.92 84.59
WePe top-10 89.57 86.67 82.62 79.33 78.95 74.42 77.15 72.29 92.65 91.36 91.91 90.60 93.77 92.71 93.17 91.76 88.42 86.46 87.58 85.07
WePe top-12 89.23 87.86 84.38 81.19 78.63 74.13 75.33 70.50 94.29 93.81 92.53 91.71 94.64 94.32 93.15 92.15 89.90 88.22 88.01 85.99
WePe top-14 89.69 87.57 82.60 79.24 79.69 76.06 76.74 71.26 93.05 92.30 92.45 91.23 94.71 94.78 94.96 94.22 89.44 88.14 88.15 86.09
WePe top-16 90.58 89.40 84.80 82.08 80.28 76.54 76.57 72.88 92.81 92.55 92.11 91.10 92.89 92.72 93.05 92.26 91.46 90.60 88.28 86.68
WePe top-18 90.02 87.83 83.39 80.58 79.12 74.64 76.18 71.12 91.82 91.36 92.26 91.71 93.77 93.39 93.68 92.89 89.12 87.57 87.71 85.68
A.15 Details of Datasets
IMAGENET. The real images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval. The images are provided by (Stein et al., 2023). The generative model includes: ADM, ADMG, BigGAN, DiT-XL-2, GigaGAN, LDM, StyleGAN-XL, RQ-Transformer and Mask-GIT. The resolution of real images and generated images are $256\times 256$. We crop the image randomly to $224\times 224$ resolution.
LSUN-BEDROOM. The real images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval. The images are provided by (Stein et al., 2023). The generative model includes: ADM, DDPM, iDDPM, StyleGAN, Diffusion-Projected GAN, Projected GAN and Unleashing Transformers. The resolution of real images and generated images are $256\times 256$. We crop the image randomly to $224\times 224$ resolution.
GenImagThe real images and generated images can be obtained at https://github.com/GenImage-Dataset/GenImage. The images are provided by (Zhu et al., 2023). The generative model includes: Midjourney, SD V1.4, SD V1.5, ADM, GLIDE, Wukong, VQDM and BigGAN. The real images come from ImageNet, and different images have different resolutions. Following (Stein et al., 2023), we resize the image to $256\times 256$ resolution and adjust its format to keep the same with the generated images, then we randomly crop it to $224\times 224$ resolution to extract features. For the generated images, we report our processing in detail as follows:
  • Midjourney. The resolution of images generated by Midjourney is $1024\times 1024$, and we randomly crop them to $224\times 224$ resolution.
  • SD V1.4. The resolution of images generated by SD V1.4 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
  • SD V1.5. The resolution of images generated by SD V1.5 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
  • ADM. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
  • GLIDE. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
  • Wukong. The resolution of images generated by SD V1.5 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
  • VQDM. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
  • BigGAN. The resolution of images generated by SD V1.5 is $128\times 128$, and we fill them with zero pixels to $224\times 224$ resolution.

### v2

@@
-Detecting Discrepancies Between AI-Generated and Natural Images Using Uncertainty
+Epistemic Uncertainty for Generated Image Detection
 Abstract
-In this work, we propose a novel approach for detecting AI-generated images by leveraging predictive uncertainty to mitigate misuse and associated risks.
-The motivation arises from the fundamental assumption regarding the distributional discrepancy between natural and AI-generated images.
-The feasibility of distinguishing natural images from AI-generated ones is grounded in the distribution discrepancy between them.
-Predictive uncertainty offers an effective approach for capturing distribution shifts, thereby providing insights into detecting AI-generated images.
-Namely, as the distribution shift between training and testing data increases, model performance typically degrades, often accompanied by increased predictive uncertainty.
-Therefore, we propose to employ predictive uncertainty to reflect the discrepancies between AI-generated and natural images.
-In this context, the challenge lies in ensuring that the model has been trained over sufficient natural images to avoid the risk of determining the distribution of natural images as that of generated images.
-We propose to leverage large-scale pre-trained models to calculate the uncertainty as the score for detecting AI-generated images.
-This leads to a simple yet effective method for detecting AI-generated images using large-scale vision models: images that induce high uncertainty are identified as AI-generated.
-Comprehensive experiments across multiple benchmarks demonstrate the effectiveness of our method.
+We introduce a novel framework for AI-generated image detection through epistemic uncertainty, aiming to address critical security concerns in the era of generative models.
+Our key insight stems from the observation that distributional discrepancies between training and testing data manifest distinctively in the epistemic uncertainty space of machine learning models.
+In this context, the distribution shift between natural and generated images leads to elevated epistemic uncertainty in models trained on natural images when evaluating generated ones.
+Hence, we exploit this phenomenon by using epistemic uncertainty as a proxy for detecting generated images.
+This converts the challenge of generated image detection into the problem of uncertainty estimation, underscoring the generalization performance of the model used for uncertainty estimation.
+Fortunately, advanced large-scale vision models pre-trained on extensive natural images have shown excellent generalization performance for various scenarios.
+Thus, we utilize these pre-trained models to estimate the epistemic uncertainty of images and flag those with high uncertainty as generated.
+Extensive experiments demonstrate the efficacy of our method.
+Code is available at https://github.com/tmlr-group/WePe.
 1 Introduction
-Recent advancements in generative models have revolutionized image generation, enabling the production of highly realistic images (Midjourney, ; Wukong, ; Rombach et al., 2022).
+Recent advancements in generative models have revolutionized image generation, enabling the production of highly realistic images (Midjourney,, 2022; Wukong,, 2022; Rombach et al.,, 2022).
 Despite the remarkable capabilities of these models, they pose significant challenges, particularly the rise of deepfakes and manipulated content.
@@
 In response to these critical concerns, a variety of methodologies for detecting generated images have emerged.
-A prevalent strategy treats this detection task as a binary classification problem, necessitating the collection of extensive datasets comprising both natural and AI-generated images to train classifiers (Wang et al., 2020).
-While existing detection methods have demonstrated notable successes, they typically encounter challenges in generalizing to images produced by previously unseen generative models Wang et al.
-(2023a).
-One promising avenue to enhance the robustness of detection capabilities involves constructing more extensive training datasets by accumulating a diverse array of natural and synthetic images.
-However, these attempts are often computationally intensive, requiring substantial datasets for effective binary classification.
-Additionally, maintaining robust detection necessitates continually acquiring images generated by the latest models.
+A prevalent strategy treats this task as a binary classification problem, necessitating the collection of extensive datasets comprising both natural and AI-generated images to train classifiers (Wang et al.,, 2020).
+While existing methods have demonstrated notable successes, unseen generative models Wang et al., (2023) pose challenges for them in generalizing to images with distribution shifts.
+One promising avenue to enhance the robustness of detection capabilities involves constructing more extensive training sets by accumulating a diverse array of natural and generated images.
+However, these attempts are computationally intensive, requiring substantial datasets for effective classification.
+Besides, maintaining robust detection necessitates continually acquiring images generated by the latest generative models.
 And when the latest generative models are not open-sourced, acquiring a large number of generated images to train classifiers is challenging.
-This highlights the urgent need for a novel framework to detect AI-generated images without requiring AI-generated images.
-Figure 1: Large-scale fundamental models trained on a large number of real images are capable of distinguishing between real and generated images.
-A recent work (Tan et al., 2024) shows that features extracted by ViT of CLIP (Radford et al., 2021) can be employed to separate natural and AI-generated images, motivating an effective approach to detecting images by training a binary classifier in the feature space of CLIP.
-This provides a promising direction to explore the possibility that large-scale foundational models already have the ability to capture the subtle differences between real images and AI-generated images.
-This may be similar to the emergence phenomenon in large language models (Wei et al., 2022), even if these models are not designed to distinguish between real and generated images.
-This phenomenon may arise from the training data of large vision models.
-In particular, since training models with AI-generated data usually leads to model collapse (Shumailov et al., 2024), current large vision models are mainly trained over natural data, making these models biased between natural and AI-generated data.
-Even though natural and AI-generated images appear extremely semantically similar to humans, such “biased” large models can capture the difference between natural and generated images.
-The distribution discrepancy in features between natural and generated images motivates us to revisit the strategy of detecting AI-generated images.
-Specifically, the ability of humans to distinguish between natural and generated images relies on the existence of the discrepancy between them.
-Moreover, humans are not trained to distinguish natural and generated images.
-To verify this intuition, we visualize the features of natural and generated images using a large-scale visual fundamental model (DINOv2), following previous work (Tan et al., 2024).
+This highlights the urgent need for a novel framework to detect generated images without reliance on generated images.
+Figure 1: Models trained on a large number of natural images are capable of distinguishing between natural and generated images.
+A recent work (Tan et al.,, 2024) shows that features extracted by ViT of CLIP (Radford et al.,, 2021) can be employed to separate natural and AI-generated images, motivating an effective approach to detecting images by training a binary classifier in the feature space of CLIP.
+This provides a promising direction to explore the possibility that large-scale foundational models already have the ability to capture the subtle differences between natural images and AI-generated images.
 As shown in Figure 1, even for images sampled from the same class, there are large distributional discrepancies in the feature space of DINOv2.
-Building on this foundation, we propose to leverage predictive uncertainty as the score for distinguishing between natural and AI-generated images.
-This is because predictive uncertainty offers an effective approach for capturing distribution shifts.
-In particular, recent studies (Snoek et al., 2019; Schwaiger et al., 2020) indicate that models tend to show increased uncertainty for out-of-distribution (OOD) samples.
-Hence, for large vision models trained only on natural images, we can treat the natural images as in-distribution samples and the generated images as OOD samples.
+This observation is consistent with an important metric for evaluating generative models-the FID score.
+FID score measures feature distribution discrepancy between natural and generated images on the Inception network (Szegedy et al.,, 2015).
+A FID score of 0 indicates that there is no difference between the two distributions.
+However, even on these simple networks such as Inception v3, advanced generative models like ADM still achieve an FID score of 11.84, not to mention that on powerful models such as DINOv2, we observe significant feature distribution discrepancy.
+However, despite these distributional differences, the inherent diversity of natural images precludes direct modeling of their distribution.
+This challenge motivates us to consider an alternative approach to reflect the distributional differences between natural and generated images.
+Figure 2: WePe reflects the distribution discrepancy between AI-generated and natural images.
+In this study, we exploit the distributional disparity between natural and generated images by leveraging epistemic uncertainty to differentiate them.
+Epistemic uncertainty quantifies a model’s confidence in its predictions, reflecting its knowledge of the data distribution.
+As the volume of training data increases, a model’s epistemic uncertainty for in-distribution (ID) data diminishes.
+For a foundational model pre-trained on an extensive dataset of natural images, we posit that its epistemic uncertainty is lower for natural images compared to generated ones, due to the model’s alignment with the natural distribution.
+This perception is consistent with recent studies (Snoek et al.,, 2019; Schwaiger et al.,, 2020), which indicate that models tend to show increased uncertainty for out-of-distribution (OOD) samples.
 The challenge comes from efficiently obtaining the uncertainty of the model on the test samples.
-Classical approaches include Monte-Carlo Dropout (MC-Dropout) (Gal & Ghahramani, 2016) and Deep Ensembles (Lakshminarayanan et al., 2017).
-However, in our attempts, MC-Dropout obtains sub-optimal results (in Table 5), and it is not practical to train multiple large models independently for ensemble.
-Instead, we find that this uncertainty can be well captured by perturbing the weights of the models.
-As shown in Figure 2, when a moderate level of perturbation is applied, the real image has consistent features on the model before and after the perturbation, but the generated image has large differences in features on the model before and after the perturbation.
+Classical approaches include Monte-Carlo Dropout (MC-Dropout) (Gal and Ghahramani,, 2016) and Deep Ensembles (Lakshminarayanan et al.,, 2017).
+However, in our attempts, MC-Dropout obtains sub-optimal results (in Table 2), and it is challenging to train multiple large models independently for model ensemble.
+Instead, our theoretical results show that this uncertainty can be well captured by perturbing the weights of the models.
+As shown in Figure 3, when a moderate level of perturbation is applied, the natural image has consistent features on the model before and after the perturbation, but the generated image has large differences in features on the model before and after the perturbation.
 In this paper, we propose a novel method for AI-generated image detection by weight perturbation (WePe).
 Our hypothesis is that the model has greater uncertainty in predicting the OOD sample compared to the ID sample, and that this uncertainty can be expressed through sensitivity to weight perturbations.
-For a large model trained on a large number of real images, the real images can be considered ID samples, while the generated images are considered OOD samples.
-Thus, the sensitivity of the samples to the weight perturbation of the large model can be an important indicator to determine whether the generative models generate the samples.
-Despite its simplicity, WePe achieves state-of-the-art performance on various benchmarks.
-Figure 2: Natural and generated images are differently sensitive to perturbations in the model weights.
-When a moderate perturbation is applied ($0.1$), the natural image has essentially the same features on the model before and after the perturbation, but the generated image already shows a large difference.
+For a large model trained on a large number of natural images, the natural images can be considered ID samples, while the generated images are considered OOD samples.
+Thus, the sensitivity of the samples to the weight perturbation of the large model can be an important indicator to determine whether the sample is generated by the generative models or not.
+As shown in Figure 2, we calculate FID scores between different types of generated images and natural images on DINOv2.
+It shows the performance of WePe strongly correlates with FID score, indicating the effectiveness of WePe in detecting distribution discrepancy.
+Figure 3: Natural and generated images exhibit distinct sensitivities to perturbations in model weights.
+A moderate perturbation ($0.1$) results in minimal changes to the features of the natural image, while the generated image shows significant differences.
 We summarize our main contributions as follows:
   • We provide a new perspective to detect AI-generated images by calculating predictive uncertainty.
-This is built upon an intuitive assumption that the natural and generated images differ in distribution, making it possible to employ uncertainty to represent the distribution discrepancy.
-  • We propose to leverage large vision models to calculate the predictive uncertainty.
+This is built on the fact that natural and generated images differ in data distribution, making it possible to employ uncertainty to represent the distribution discrepancy.
+  • Since the data distribution discrepancy between generated and natural images is reflected in feature distribution discrepancy on the vision model, we propose using a large vision model to compute prediction uncertainty to highlight this difference.
 The intuition is that large vision models are merely trained on natural images, making it possible to exhibit different uncertainties about natural and generated images.
-We capture this uncertainty by weight perturbation to effectively detect images (Eq.
-3).
-  • Comprehensive experiments on multiple benchmarks demonstrate that the proposed method outperforms previous methods, including training-based methods, achieving state-of-the-art performance.
-2 Related Works
-AI-Generated images detection.
-Recent advancements in generative models, such as those by (Brock et al., 2019; Ho et al., 2020), have led to the creation of highly realistic images, highlighting the urgent need for effective algorithms to distinguish between natural and generated images.
-Prior research, including works by (Frank et al., 2020; Marra et al., 2018), primarily focuses on developing specialized binary classification neural networks to differentiate between natural and generated images.
-Notably, CNNspot (Wang et al., 2020) demonstrates that a standard image classifier trained on ProGAN can generalize across various architectures when combined with specific data augmentation techniques.
-F3Net (Qian et al., 2020) distinguishes real face images from fake face images with the help of frequency statistical differences.
-NPR (Tan et al., 2024) introduces the concept of neighboring pixel relationships to capture differences between natural and generated images.
-Although these methods show superior performance on generators in the training set, they often do not generalize well to unknown generators.
-In addition to this, training-based methods are susceptible to small perturbations in the image.
-For this reason, recently, some training-free methods have been proposed.
-AEROBLADE (Ricker et al., 2024) calculates the reconstruction error with the help of the autoencoder used in latent diffusion models (Rombach et al., 2022).
-RIGID (He et al., 2024) finds that real images are more robust to small noise perturbations than generated images in the representation space of the vision foundation models and exploits this property for detection.
-However, these methods usually make overly strong assumptions about natural or generated images, leading to insufficient generalization.
-In our paper, we propose a training-free detection method through uncertainty analysis.
-Based on the widespread phenomenon that generated images have greater uncertainty than real images on models trained with real images, our method achieves robust detection performance.
-Uncertainty estimation.
-Uncertainty estimation in machine learning has seen significant advancements in recent years.
-(Gal & Ghahramani, 2016) introduces Monte Carlo Dropout (MC Dropout), which uses dropout at inference to estimate uncertainty from the variance of multiple predictions.
-(Lakshminarayanan et al., 2017) develops deep ensembles, demonstrating improved uncertainty estimates through training multiple model indepently with different initializations.
-Recent work by (Snoek et al., 2019) analyzes the calibration of uncertainty in deep learning models, highlighting the importance of reliable uncertainty measures.
-Additionally, (Guo et al., 2017) explore the use of temperature scaling to enhance the calibration of model predictions.
-In our paper, We measure uncertainty by perturbing the model’s weights.
-Random weight perturbations.
-Weight perturbation, i.e.
-adding noise to network weights, has been extensively studied.
-Many methods (Khan et al., 2018; Wu et al., 2020) perform training regularization by adding noise to the weights of the neural network during training to improve the generalization of the network.
-Weight perturbation has also been applied in adversarial attacks to study the robustness of the network (He et al., 2019; Garg et al., 2020).
-In addition to this, some works also study the sensitivity of neural network to weights perturbation.
-(Cheney et al., 2017) shows that convolutional networks are surprisingly robust to a number of internal perturbations in the higher convolutional layers but the bottom convolutional layers are much more fragile.
-(Weng et al., 2020) proposes an efficient approach to compute a certified robustness bound of weight perturbations.
-In our work, we find neural networks trained on natural images exhibit different robustness to weight perturbations for natural and generated images.
-Based on this property, we propose an efficient algorithm to distinguish between natural and generated images.
-3 Method
+We capture this uncertainty through weight perturbation, enabling effective detection.
+(Eq.
+13).
+  • Extensive experiments across multiple benchmarks show the proposed method surpasses existing methods, highlighting the efficacy of uncertainty in detecting AI-generated images.
+2 Preliminaries
+2.1 AI-generated image detection
+The task of AI-generated image detection involves classifying a test image $\mathbf{x}$ as either a natural image or one produced by a generative model.
+Traditional supervised learning approaches rely on a curated dataset comprising both natural and AI-generated images to train a feature extractor and a binary classifier.
+This process is formalized as follows.
+Let $\mathcal{D}^{0}=\{\mathbf{x}^{0}_{1},\mathbf{x}^{0}_{2},\dots,\mathbf{x}^{0}_{N^{0}}\}$ denote a set of $N^{0}$ AI-generated images, each labeled as $y=0$ (generated), and $\mathcal{D}^{1}=\{\mathbf{x}^{1}_{1},\mathbf{x}^{1}_{2},\dots,\mathbf{x}^{1}_{N^{1}}\}$ denote a set of $N^{1}$ natural images, each labeled as $y=1$ (natural).
+The objective is to learn a feature extractor $F(\cdot;\theta_{F})$, parameterized by $\theta_{F}$, and a binary classifier $D(\cdot;\theta_{D})$, parameterized by $\theta_{D}$, by minimizing a loss function $\ell(\cdot)$ over the training data:
+$F,D=\arg\min_{\theta_{F},\theta_{D}}\ell\left(D\left(F(\mathbf{x};\theta_{F});\theta_{D}\right),y\right),$ (1)
+where $y\in\{0,1\}$ represents the ground-truth label of the input image $\mathbf{x}$.
+Upon completion of training, the feature extractor and classifier are used to compute a decision score for a test image $\mathbf{x}$, defined as $S(\mathbf{x})=D(F(\mathbf{x};\theta_{F});\theta_{D})$.
+The source of the image is determined by comparing this score to a predefined threshold $\tau$:
+$\operatorname{pred}(\mathbf{x})=\begin{cases}\text{generated},&\text{if }S(\mathbf{x})<\tau,\\ \text{natural},&\text{otherwise.}\end{cases}$ (2)
+The robustness and generalization of such methods depend critically on the size and diversity of the training dataset.
+To address this, techniques such as CNNspot employ sophisticated data augmentation strategies, including Gaussian blur and JPEG compression, to enhance the variability of the training data.
+In contrast, UnivFD adopts a different approach by utilizing the CLIP model as a fixed feature extractor and training only a single linear classification layer.
+Despite these innovations, both methods exhibit limited generalization when applied to images produced by unseen generative models.
+2.2 Bayesian neural networks and uncertainty estimation
+Traditional neural networks map inputs $\mathbf{x}$ to outputs $\mathbf{y}$ through a parameterized function $f(\mathbf{x};\theta)$, where weights $\theta$ are optimized deterministically, offering no uncertainty quantification.
+Bayesian Neural Networks (Neal,, 2012) address this by modeling $\theta$ as random variables with a prior distribution $p(\theta)$.
+Given a dataset $\mathcal{D}$, the posterior is computed via Bayes’ theorem:
+$p(\theta|\mathcal{D})=\frac{p(\mathcal{D}|\theta)p(\theta)}{p(\mathcal{D})}.$ (3)
+For a new input $\mathbf{x}^{*}$, the posterior predictive distribution:
+$p(\mathbf{y}^{*}|\mathbf{x}^{*},\mathcal{D})=\int\underbrace{p(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)}_{\text{Aleatoric}}\underbrace{p(\theta|\mathcal{D})}_{\text{Epistemic}}d\theta$ (4)
+captures prediction uncertainty.
+Exact posterior inference is intractable due to the high dimensionality of $\theta$.
+Monte Carlo methods, such as MC Dropout (Gal and Ghahramani,, 2016), approximate the posterior by using dropout during both training and testing and generating multiple stochastic predictions.
+Uncertainty is quantified through the variance or entropy of these predictions, reflecting two types:
+  • Epistemic Uncertainty: Uncertainty in model parameters, encoded in $p(\mathbf{w}|\mathcal{D})$, which decreases with more data.
+  • Aleatoric Uncertainty: Inherent noise in the data, modeled by the likelihood $p(\mathbf{y}|\mathbf{x},\mathbf{w})$.
+3 Uncertainty based AI-generated image detection
 3.1 Motivation
-Our method is built upon a foundational assumption that natural and generated images have different distributions.
-This is a reasonable assumption; otherwise, we cannot distinguish between natural and generated images.
-Fortunately, this assumption is consistent with previous work (Tan et al., 2024) and our empirical observations, as shown in Figure 1.
-Hence, for the large models trained merely on natural images, we can regard natural images as in-distribution (ID) data while generated images as out-of-distribution (OOD) data.
-This distribution discrepancy can be reflected by the widely used predictive uncertainty, since neural networks typically exhibit higher uncertainty for OOD samples (Snoek et al., 2019; Schwaiger et al., 2020).
-This leads to a simple yet novel approach to determine whether a test image is generated by AI models when we can calculate its uncertainty on a pre-trained large vision model.
-In the following, we first give a complete introduction to our method WePe, then describe on our reasons for choosing the method, and finally, we explore why WePe works.
-3.2 Uncertainty Definition
-Classical methods of uncertainty estimation, such as Deep Ensembles and MC Dropout, can simply be viewed as using the variance of the results of multiple predictions as an estimation of uncertainty $u(\mathbf{x})$:
-$\mu(\mathbf{x})=\frac{1}{n}\sum_{t=1}^{n}\hat{y}_{t}(\mathbf{x}),\quad u(\mathbf{x})=\sigma^{2}=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{y}_{t}(\mathbf{x})-\mu(\mathbf{x})\right)^{2},$ (1)
+Uncertainty estimation is a well-established field critical to deep learning practitioners, as it facilitates explicit handling of uncertain inputs and edge cases (Durasov et al.,, 2021; Everett et al.,, 2022).
+By quantifying model confidence, practitioners can make informed decisions, such as deferring high-uncertainty inputs to human evaluation in classification tasks, thereby improving model reliability and robustness.
+In this study, we do not propose a novel method for uncertainty estimation.
+Instead, we focus on distinguishing natural from generated images by leveraging epistemic uncertainty as a discriminative metric.
+Our approach is grounded in the well-established principle that (Lahlou et al.,, 2023; Gal and Ghahramani,, 2016):
+Epistemic uncertainty, which reflects a model’s lack of knowledge, can be reduced by acquiring additional information.
+Formally, we adopt a Bayesian framework to quantify epistemic uncertainty through the posterior distribution over model parameters:
+$p(\theta|\mathcal{D})=\frac{p(\mathcal{D}|\theta)p(\theta)}{p(\mathcal{D})},$ (5)
+where $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denotes the training dataset, with samples drawn i.i.d.
+from $p(\mathbf{y}|\mathbf{x},\theta_{0})$.
+Epistemic uncertainty is captured by the posterior variance:
+$\text{Var}(\theta|\mathcal{D})=\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}[(\theta-\mathbb{E}[\theta|\mathcal{D}])^{2}].$ (6)
+Under regularity conditions, the Bernstein-von Mises theorem (Van der Vaart,, 2000) establishes that, as the sample size $N\to\infty$,
+$p(\theta|\mathcal{D})\approx\mathcal{N}\left(\hat{\theta}_{N},\frac{1}{N}I(\theta_{0})^{-1}\right),$ (7)
+where $\hat{\theta}_{N}$ is the maximum likelihood estimate, and $I(\theta_{0})=\mathbb{E}\left[\left(\frac{\partial\log p(\mathbf{y}|\mathbf{x},\theta)}{\partial\theta}\right)^{2}\big|_{\theta=\theta_{0}}\right]$ represents the Fisher information matrix.
+Here, $\theta_{0}$ approximates the true parameters of the natural image distribution, as learned by a pre-trained vision model.
+Consequently, $\text{Var}(\theta|\mathcal{D})\propto\frac{1}{N}$, indicating that epistemic uncertainty decreases as the training data volume increases.
+For a test input $\mathbf{x}^{*}$ drawn from a distribution $p_{\text{test}}(\mathbf{x})\neq p_{\text{train}}(\mathbf{x})$, the predictive distribution is given by:
+$p(\mathbf{y}^{*}|\mathbf{x}^{*},\mathcal{D})=\int p(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)p(\theta|\mathcal{D})\,d\theta.$ (8)
+This distribution exhibits elevated variance due to the misalignment between $p(\theta|\mathcal{D})$ and the parameters required to model $p_{\text{test}}(\mathbf{x})$.
+The predictive variance can be decomposed as:
+$\text{Var}(\mathbf{y}^{*}|\mathbf{x}^{*},\mathcal{D})=\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}[\text{Var}(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)]+\text{Var}_{\theta\sim p(\theta|\mathcal{D})}[\mathbb{E}(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)],$ (9)
+where the second term, $\text{Var}_{\theta\sim p(\theta|\mathcal{D})}[\mathbb{E}(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)]$, corresponds to epistemic uncertainty and dominates for out-of-distribution (OOD) samples.
+Under distribution shift, the Fisher information $I(\theta_{0})$, defined with respect to $p_{\text{train}}(\mathbf{x})$, is misaligned with $p_{\text{test}}(\mathbf{x})$, limiting the reduction of posterior variance and resulting in persistently high epistemic uncertainty (Snoek et al.,, 2019).
+Based on the above analysis, as well as the phenomenon of distributional shifts between natural and generated images on the foundational model (see Figures 1 and 8), we believe that the difference in the epistemic uncertainty of the foundational model on natural and generated images can be a valid indicator to distinguish between them.
+3.2 Uncertainty estimation via weight perturbation
+Classical methods of epistemic uncertainty estimation, such as Deep Ensembles and MC Dropout, can simply be viewed as using variance of multiple prediction results as an estimation of uncertainty $u(\mathbf{x})$:
+$\mu(\mathbf{x})=\frac{1}{n}\sum_{t=1}^{n}\hat{y}_{t}(\mathbf{x}),u(\mathbf{x})=\sigma^{2}=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{y}_{t}(\mathbf{x})-\mu(\mathbf{x})\right)^{2},$ (10)
 where, $\hat{y}_{t}$ denotes the $t$-th prediction.
 The multiple predictions of Deep Ensembles come from multiple independently trained neural networks, while the multiple predictions of MC Dropout come from the use of dropout during inference, which can be regarded as multiple prediction using neural networks with different structures.
-3.3 Uncertainty Calculation
-The predictive uncertainty is typically calculated as the variance of predictions obtained with certain perturbations.
-In this work, we simply leverage $\theta$ as the features or parameters before perturbation.
-Specifically, the predictive uncertainty $u(\mathbf{x})$ can be calculated by,
+In this study, we leverage DINOv2 for uncertainty estimation, capitalizing on two key advantages.
+First, DINOv2 is pre-trained on an extensive datasets, endowing it with robust generalization capabilities.
+This broad pre-training enables the model to effectively capture the distributional characteristics of natural images, thereby enhancing the generalizability of our proposed method across diverse scenarios.
+Second, as DINOv2 is exclusively pre-trained on natural images, it exhibits distinct uncertainty profiles when processing natural versus AI-generated images.
+This differential uncertainty provides a reliable basis for distinguishing between the two image types.
+DINOv2 is a self-supervised learning model, employing a student-teacher framework where the student model $\theta$ is trained to align with the representations of the teacher model $\theta_{t}$.
+Consequently, we utilize the feature similarity between the embeddings produced by the teacher and student models as the prediction:
+$\hat{y}(x)=f(\mathbf{x};\theta)^{\top}f(\mathbf{x};\theta_{t}),$ (11)
+where $f(\mathbf{x};\theta)$ denotes the L2-normalized features of an input image $\mathbf{x}$ when inferring with the parameter $\theta$.
+However, estimating uncertainty with DINOv2 presents challenges in obtaining multiple predictions.
+First, as DINOv2 does not employ dropout during training, MC Dropout may yield suboptimal results (see Table 2).
+Second, training multiple DINOv2-scale models is computationally infeasible, rendering Deep Ensemble impractical.
+Instead, we use an alternative approach, weight perturbation, to obtain multiple predictions.
+Specifically, let $\theta_{k}$ denote the perturbed parameters of the student network, according to Eq.
+10, the predictive uncertainty $u(\mathbf{x})$ can be calculated by,
 $u(\mathbf{x})=\frac{1}{n}\sum_{k=1}^{n}[f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta_{t})-\sum_{j=1}^{n}\frac{f(\mathbf{x};\theta_{j})^{\top}f(\mathbf{x};\theta_{t})}{n}]^{2},$ (12)
-where $n$ is the number of samples, $f(\mathbf{x};\theta_{k})$ denotes the L2-normalized features of an input image $\mathbf{x}$ when inferring with the parameter $\theta_{k}$, and $\theta_{t}$ stands for the teacher model used in DINOv2.
 However, we cannot access the teacher model $\theta_{t}$, making it challenging to calculate the uncertainty.
@@
 This can be formalized by,
-$u(\mathbf{x})\leq\frac{1}{n}\sum_{k=1}^{n}\left\|f(\mathbf{x};\theta_{k})-\frac{1}{n}\sum_{j}^{n}f(\mathbf{x};\theta_{j})\right\|^{2}\left\|f(\mathbf{x};\theta_{t})\right\|^{2}=2-\frac{2}{n}\sum_{k=1}^{n}f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta),$ (3)
-where $\theta$ denotes the parameter before injecting perturbation, and we leverage an unbiased assumption that the expectation $\mathbb{E}_{\theta_{j}}f(\mathbf{x};\theta_{j})$ approaches the feature $f(\mathbf{x};\theta)$ extracted by the non-perturbed parameter $\theta$.
+$\begin{split}u(\mathbf{x})\leq\frac{1}{n}\sum_{k=1}^{n}\left\|f(\mathbf{x};\theta_{k})-\frac{1}{n}\sum_{j}^{n}f(\mathbf{x};\theta_{j})\right\|^{2}\left\|f(\mathbf{x};\theta_{t})\right\|^{2}=2-\frac{2}{n}\sum_{k=1}^{n}f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta),\end{split}$ (13)
+where $\theta$ denotes the parameter of student model before injecting perturbation, and we leverage an unbiased assumption that the expectation $\mathbb{E}_{\theta_{j}}f(\mathbf{x};\theta_{j})$ approaches the feature $f(\mathbf{x};\theta)$ extracted by the non-perturbed parameter $\theta$.
 Eq.
 13 provides a simple approach to calculate the uncertainty without needing a teacher model used in the training phase of DINOv2.
+We provide an analysis of the validity of this upper bound in Appendix D.
 The insight of Eq.
@@
 Specifically, if an image $\mathbf{x}$ causes a high feature similarity between the original and perturbed parameter, the image leads to a low uncertainty and is more likely to be a natural image.
-3.4 An overview of WePe
-As discussed above, the proposed WePe is based on a large model pre-trained on a large number of natural images.
-In this work, we chose DINOv2 (Oquab et al., 2024), a large model trained with contrastive learning on image data.
-In order to capture the uncertainty of the model on the test images, we extract image features using the original model and the model after adding noise to the parameters respectively.
-The similarity between the pre-perturbation and post-perturbation feature vectors is quantified using a suitable distance metric, such as cosine similarity.
-Images exhibiting high similarity are classified as real, while those with low similarity are identified as generated.
-This method not only capitalizes on the characteristics of the DINOv2 model but also provides a robust framework for distinguishing between real and generated images based on their feature stability under model perturbation.
-3.5 Discussions
-Why choose DINOv2?
-In addition to DINOv2, CLIP is a commonly used model.
-However, our experiments show that CLIP performs sub-optimally compared to DINOv2 (see Table 5).
-We believe this difference stems from their training strategies.
-Unlike DINOv2, CLIP is a multimodal model that combines image and textual features from captions for contrastive learning, which may lead it to focus on broader semantic features rather than fine details.
-In contrast, DINOv2 emphasizes contrastive learning solely on images, allowing it to better capture subtle differences between natural and generated images.
-Therefore, we use DINOv2 in our experiments.
-Why choose weight perturbation?
-Common methods for measuring uncertainty include MC Dropout and Deep Ensembles.
-MC Dropout involves keeping dropout active during testing and performing multiple forward passes on the inputs to generate outputs with different network structures.
-The variability among these outputs serves as an estimate of the model’s uncertainty regarding the input data.
-However, since DINOv2 does not utilize dropout during training, MC Dropout may not yield optimal results (see Table 5).
-Deep Ensembles, on the other hand, trains multiple networks independently and uses the differences in their outputs on test samples to assess uncertainty.
-However, training multiple DINOv2-level models is impractical.
-Therefore, in our study, we choose to perturb the model parameters and assess the differences in outputs from the original and perturbed models to estimate uncertainty for the test images.
-Why does weight perturbation work?
-Incorporating weight perturbations during testing can effectively simulate Bayesian inference, thereby capturing the inherent uncertainty in neural networks.
-From a Bayesian perspective, the weights of a neural network are not fixed but rather distributions reflecting the range of plausible values given the data.
-By introducing noise into the network weights at test time, we mimic the process of drawing samples from a posterior distribution over weights, a core concept in Bayesian inference.
-This technique enables the model to generate diverse predictions, reflecting its uncertainty, particularly when encountering out-of-distribution samples.
-This approach is akin to Bayesian neural networks, where weight uncertainty is explicitly modeled.
-For example, (Blundell et al., 2015) proposed using variational methods to approximate weight distributions in Bayesian neural networks, allowing for uncertainty quantification through weight perturbations.
-Similarly, (Gal & Ghahramani, 2016) demonstrated that introducing dropout at test time serves as a Bayesian approximation, with the added noise acting as a proxy for weight sampling, thus allowing for reliable uncertainty estimation.
-In this paper, we capture uncertainty by adding noise directly to the model parameters to distinguish between natural and generated images.
-However, trying to theoretically analyze the different sensitivities of natural and generated images to model weights in the representation space of a large model is a difficult task.
-Therefore, we only draw this conclusion through various empirical observations in this paper, and do not directly prove this theoretically.
-The theoretical proof will be left to our future work.
-4 Experiments
-4.1 Experiment setup
-Datasets.
-Following previous work (He et al., 2024), we evaluate the performance of WePe on ImageNet (Deng et al., 2009), LSUN-BEDROOM (Yu et al., 2015) and GenImage (Zhu et al., 2023).
-For ImageNet and LSUN-BEDROOM, the generated images are provided by  (Stein et al., 2023).
-For ImageNet, the generative models include ADM (Dhariwal & Nichol, 2021), ADM-G, LDM (Rombach et al., 2022), DiT-XL2 (Peebles & Xie, 2023), BigGAN (Brock et al., 2019), GigaGAN (Kang et al., 2023), StyleGAN (Karras et al., 2019), RQ-Transformer (Lee et al., 2022), and MaskGIT (Chang et al., 2022).
-For LSUN-BEDROOM,the generative models include ADM, DDPM (Ho et al., 2020), iDDPM (Nichol & Dhariwal, 2021), Diffusion Projected GAN (Wang et al., 2023b), Projected GAN (Wang et al., 2023b), StyleGAN (Karras et al., 2019) and Unleasing Transformer (Bond-Taylor et al., 2022).
-GenImage primarily employs the Diffusion model for image generation.
-The fake images are generated by Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), GLIDE, VQDM (Gu et al., 2022), Wukong (Wukong, ), BigGAN, ADM, and Midjourney (Midjourney, ).
-It contains 1,331,167 real and 1,350,000 generated images.
-Evaluation metrics.
-Following RIGID, we mainly use the following metrics: (1) the average precision (AP), (2) the area under the receiver operating characteristic curve (AUROC).
-On GenImage, since the dataset is too large and re-implementing all baselines is time-consuming, we obtain the results directly from the corresponding papers and report the classification accuracy (ACC).
-Baselines.
-Following RIGID, we take both training-free methods and training methods as baselines.
-For training-free methods, we take RIGID (He et al., 2024) and AEROBLADE (Ricker et al., 2024) as baselines.
-For training methods, we take DIRE (Wang et al., 2023a), CNNspot (Wang et al., 2020), Ojha (Ojha et al., 2023) and NPR (Tan et al., 2024) as baselines.
-Besides, on GenImage, we also report the result of Frank (Frank et al., 2020), Durall (Durall et al., 2020), Patchfor (Chai et al., 2020), F3Net (Qian et al., 2020), SelfBland (Shiohara & Yamasaki, 2022), GANDetection (Mandelli et al., 2022), LGrad (Tan et al., 2023), ResNet-50 (He et al., 2016), DeiT-S (Touvron et al., 2021), Swin-T (Liu et al., 2021), Spec (Zhang et al., 2019), GramNet (Liu et al., 2020).
-For these baselines, we get the results directly in the corresponding papers without reimplementing them.
-Experiment details.
-To balance detection performance and efficiency, we use DINOv2 ViT-L/14.
-We report the average results under five different random seeds and report the variance in Figure 8.
-In our experiments we find that perturbing the high layers may lead to a large corruption in the features of the real images, resulting in sub-optimal results.
-Therefore, We do not perturb the high-level parameters.
-In DINOv2 ViT-L/14, the model has 24 transformer blocks, and we only perturb the parameters of the first 19 blocks with Gaussian perturbations of zero mean.
-The variance of the added Gaussian noise is proportional to the mean value of the parameters in each block, with the ratio set to 0.1.
-Ablation experiments show that our method is robust to the blocks chosen for perturbation and the level of noise.
-4.2 Results
-Comparison with other baselines.
-We conduct full comparative experiments on the three benchmarks mentioned.
-As shown in Table 1,  2 and  3, WePe achieves the best detection performance on ImageNet, LSUN-BEDROOM and GenImage without the need for training.
-It is worth noting that on the large-scale GenImage benchmark, the training-based method, despite having perfect performance on the generators used during training, performs poorly on many generators not seen during training.
-This illustrates the extreme dependence of the performance of training-based methods on the diversity of the training set.
-In contrast, our method does not require training, performs well on a wide variety of generators, and outperforms the SOTA training method by $3.36\%$ on average.
-On generators that have not been seen during training, such as VQDM, many training methods exhibit random prediction results, whereas WePe achieves superior detection performance.
-To further illustrate the effectiveness of our method, we count the difference in feature similarity between real and fake images on the pre- and post-perturbation models.
-As shown in Figure 3, the small perturbation of the model has less effect on the real images than on generated images, resulting in higher feature similarity before and after the perturbation.
-The discrepancy effectively distinguishes the real image from the generated image.
-Figure 3: Comparison of cosine similarity between original and transformed images. The generated images are from: (a) ADM, (b) BigGAN, and (c) DDPM.
-Table 1: AI-generated image detection performance on ImageNet.
-Values are percentages.
-Bold numbers are superior results and the underlined italicized values are the second-best performance.
-A higher value indicates better performance.
+Figure 4: Comparison of cosine similarity between features on original and perturbed models. The generated images are from: (a) ADM, (b) BigGAN, and (c) DDPM.
+3.3 Theoretical analysis of the effectiveness of weight perturbation
+In this section, we present a theoretical analysis elucidating why weight perturbations effectively distinguish between natural and AI-generated images.
+We establish a formal metric to quantify the sensitivity of neural network feature representations to parameter perturbations and demonstrate its differential behavior across natural images and AI-generated images.
+To formalize this notion, we introduce the following definition of perturbation sensitivity:
+Definition 3.1 (Perturbation Sensitivity).
+For a neural network $f:\mathcal{X}\to\mathbb{R}^{d}$, parameterized by $\theta\in\mathbb{R}^{p}$, which maps an input $x\in\mathcal{X}$ to a feature vector $f(x;\theta)$, the sensitivity to parameter perturbations is defined as:
+$\text{sen}(x)=\|\nabla_{\theta}f(x;\theta)\|_{F}^{2},$ (14)
+where $\nabla_{\theta}f(x;\theta)$ denotes the Jacobian matrix of the feature mapping with respect to $\theta$, and $\|\cdot\|_{F}$ represents the Frobenius norm.
+This metric captures the magnitude of variation in the feature representation induced by infinitesimal changes in the model parameters, providing a robust measure of sensitivity to weight perturbations.
+We proceed to establish the differential sensitivity of the model across distributions through the following theorem:
+Theorem 3.2 (Differential Sensitivity).
+Let a neural network $f(x;\theta)$ be trained on a large amount of natural images sampled from natural image distribution $\mathcal{D}^{1}$: $T=\{(x^{1},y^{1}),(x^{2},y^{2}),...,(x^{n},y^{n})\}\sim\mathcal{D}^{1}$.
+The expected sensitivity of the feature representations to parameter perturbations is lower for inputs drawn from $\mathcal{D}^{1}$ compared to those drawn from a generated image distribution $\mathcal{D}^{0}$, generated by generative models.
+Formally:
+$\mathbb{E}_{x\sim\mathcal{D}^{1}}\left[\text{sen}(x)\right]\leq\mathbb{E}_{x\sim\mathcal{D}^{0}}\left[\text{sen}(x)\right].$ (15)
+This theorem asserts that feature representations of natural images, optimized through training on $\mathcal{D}^{1}$, exhibit greater robustness to parameter perturbations than those of AI-generated images, which lie outside the training distribution.
+This differential sensitivity underpins the efficacy of weight perturbations in AI-generated image detection, enabling the model to distinguish natural images from AI-generated images.
+A rigorous proof of this theorem is provided in Appendix C.
+Table 1: AI-generated image detection performance on ImageNet. Values are percentages. Bold numbers are superior results. We compare training methods and training-free methods separately.
 Models
@@
 NPR 85.68 80.86 84.34 79.79 91.98 86.96 86.15 81.26 89.73 84.46 82.21 78.20 84.13 78.73 80.21 73.21 89.61 84.15 86.00 80.84
+PatchCraft 81.83 79.65 70.88 69.36 68.47 65.19 75.38 73.29 99.85 99.26 98.55 97.91 96.33 96.25 91.28 91.47 92.56 92.17 86.13 84.95
+FatFormer 91.77 90.36 83.58 83.17 92.58 92.06 86.93 85.14 98.76 98.47 97.65 98.02 97.64 97.57 96.55 95.96 97.65 97.27 93.68 93.11
+DRCT 90.26 90.07 85.74 83.85 90.24 89.88 88.27 89.06 95.87 94.99 86.89 86.12 89.11 88.39 92.38 92.41 94.44 94.47 90.36 89.92
+WePe 93.89 92.42 90.21 87.15 91.73 88.69 88.00 84.94 99.85 99.83 99.03 99.04 99.52 99.51 98.31 97.84 99.63 99.54 95.57 94.33
 Training-free Methods
-AEROBLADA 55.61 54.26 61.57 56.58 62.67 60.93 85.88 87.71 44.36 45.66 47.39 48.14 47.28 48.54 67.05 67.69 48.05 48.75 57.87 57.85
+AEROBLADE 55.61 54.26 61.57 56.58 62.67 60.93 85.88 87.71 44.36 45.66 47.39 48.14 47.28 48.54 67.05 67.69 48.05 48.75 57.87 57.85
 RIGID 87.16 85.08 80.09 77.07 72.43 69.30 70.40 65.94 90.08 89.26 86.39 84.11 86.32 85.44 90.06 88.74 89.30 89.25 83.58 81.58
 WePe 89.79 87.32 83.20 78.80 78.47 73.50 77.13 71.21 94.24 93.64 92.15 90.29 93.86 92.86 93.50 91.47 89.55 86.25 87.99 85.04
-Table 2: AI-generated image detection performance on LSUN-BEDROOM.
-Models
-ADM DDPM iDDPM Diffusion GAN Projected GAN StyleGAN Unleashing Transformer Average
-Methods AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP
-CNNspot 64.83 64.24 79.04 80.58 76.95 76.28 88.45 87.19 90.80 89.94 95.17 94.94 93.42 93.11 84.09 83.75
-Ojha 71.26 70.95 79.26 78.27 74.80 73.46 84.56 82.91 82.00 78.42 81.22 78.08 83.58 83.48 79.53 77.94
-DIRE 57.19 56.85 61.91 61.35 59.82 58.29 53.18 53.48 55.35 54.93 57.66 56.90 67.92 68.33 59.00 58.59
-NPR 75.43 72.60 91.42 90.89 89.49 88.25 76.17 74.19 75.07 74.59 68.82 63.53 84.39 83.67 80.11 78.25
-AEROBLADA 57.05 58.37 61.57 61.49 59.82 61.06 47.12 48.25 45.98 46.15 45.63 47.06 59.71 57.34 53.85 54.25
-RIGID 71.90 72.29 88.31 88.55 84.02 84.80 91.42 91.90 92.12 92.54 77.29 74.96 91.37 91.39 85.20 85.20
-WePe 73.85 70.21 88.84 87.14 86.23 83.82 94.16 93.52 95.34 95.18 83.50 80.66 94.18 93.45 88.01 86.28
-Table 3: AI-generated image detection performance on GenImage. Except for WePe and RIGID, all methods require training on SD V1.4. The results of the baseline method are all from GenImage
-Models
-Methods Midjourney SD V1.5 ADM GLIDE Wukong VQDM BigGAN Average
-ResNet-50 54.90 99.70 53.50 61.90 98.20 56.60 52.00 68.11
-DeiT-S 55.60 99.80 49.80 58.10 98.90 56.90 53.50 67.51
-Swin-T 62.10 99.80 49.80 67.60 99.10 62.30 57.60 71.19
-CNNspot 52.80 95.90 50.10 39.80 78.60 53.40 46.80 58.63
-Spec 52.00 99.20 49.70 49.80 94.80 55.60 49.80 64.41
-F3Net 50.10 99.90 49.90 50.00 99.90 49.90 49.90 64.22
-GramNet 54.20 99.10 50.30 54.60 98.90 50.80 51.70 65.66
-DIRE 60.20 99.80 50.90 55.00 99.20 50.10 50.20 66.49
-Ojha 73.20 84.00 55.20 76.90 75.60 56.90 80.30 71.73
-LaRE 66.40 87.10 66.70 81.30 85.50 84.40 74.00 77.91
-RIGID 81.54 68.72 72.35 84.15 68.57 78.98 93.02 78.19
-WePe 79.17 75.57 76.07 79.20 79.00 90.60 89.27 81.27
+Figure 5: Performance varies with perturbation intensity under different degradation mechanisms, including (a) JPEG compression, (b) Gaussian blur, and (c) Gaussian noise.
+3.4 Sharpening discriminative uncertainty through precise calibration
+The original DINOv2 model reveals notable uncertainty disparities between natural and generated images.
+With access to the training dataset, fine-tuning can amplify this uncertainty gap, enhancing detection performance.
+Specifically, we introduce the following loss function:
+$\mathcal{L}(\theta)=\mathbb{E}_{x\in X^{1}}\left[f(\mathbf{x};\theta)^{\top}f(\mathbf{x};\theta^{*})\right]-\mathbb{E}_{x\in X^{0}}\left[f(\mathbf{x};\theta)^{\top}f(\mathbf{x};\theta^{*})\right],$ (16)
+where $\theta^{*}=\theta+P$ represents a perturbed parameter set, with $P$ being a randomized perturbation matrix drawn from a predefined distribution.
+This formulation builds on the uncertainty upper bound established in Eq.
+13 to guide optimization.
+The loss encourages high confidence for natural images ($X^{1}$) and low confidence for generated images ($X^{0}$), thereby sharpening the model’s ability to distinguish between the two classes.
+We call this method WePe.
+4 Experiments
+4.1 Experiment setup
+Datasets and evaluation metrics.
+Following previous works (He et al.,, 2024; Zhu et al.,, 2023), we evaluate the performance of WePe on ImageNet (Deng et al.,, 2009), LSUN-BEDROOM (Yu et al.,, 2015), GenImage (Zhu et al.,, 2023) and DRCT-2M (Chen et al.,, 2024), with the following evaluation metrics: (1) the average precision (AP), (2) the area under the receiver operating characteristic curve (AUROC) and (3) the classification accuracy (ACC).
 Figure 6: Performance varies with variance
 Figure 7: The effects of disturbed blocks.
-Figure 6: Illustration of detection performance varying with perturbation intensity under different degradation mechanisms, (a): JPEG compression, (b): Gaussian blur, and (c): Gaussian noise.
+Baselines.
+Following RIGID, we take both training-free methods and training methods as baselines.
+For training-free methods, we take RIGID (He et al.,, 2024) and AEROBLADE (Ricker et al.,, 2024) as baselines.
+For training methods, we take DIRE (Wang et al.,, 2023), CNNspot (Wang et al.,, 2020), Ojha (Ojha et al.,, 2023), PatchCraft (Zhong et al.,, 2023), FatFormer (Liu et al.,, 2024), DRCT (Chen et al.,, 2024) and NPR (Tan et al.,, 2024) as baselines.
+Besides, on GenImage, we also report the result of Frank (Frank et al.,, 2020), Durall (Durall et al.,, 2020), Patchfor (Chai et al.,, 2020), F3Net (Qian et al.,, 2020), SelfBlend (Shiohara and Yamasaki,, 2022), GANDetection (Mandelli et al.,, 2022), LGrad (Tan et al.,, 2023), ResNet-50 (He et al.,, 2016), DeiT-S (Touvron et al.,, 2021), Swin-T (Liu et al.,, 2021), Spec (Zhang et al.,, 2019), GramNet (Liu et al.,, 2020).
+Table 2: The effect of perturbation type.
+Noise      AUROC      AP Gaussian noise      87.99      85.04 Uniform noise      89.06      86.32 Laplace noise      87.13      84.22 MC Dropout      81.63      79.71
+Experiment details.
+To balance detection performance and efficiency, we use DINOv2 ViT-L/14.
+Due to the randomness of the added noise, we report the average results under five different random seeds and report the variance in Figure 10.
+In our experiments we find that perturbing the high layers may lead to a large corruption in the features of the natural images, resulting in sub-optimal results.
+Therefore, we do not perturb the high-level parameters.
+In DINOv2 ViT-L/14, the model has 24 transformer blocks, and we only perturb the parameters of the first 19 blocks with Gaussian perturbations of zero mean.
+The variance of Gaussian noise is proportional to the mean value of the parameters in each block, with the ratio set to 0.1.
+4.2 Results
+Comparison with other baselines.
+We conduct full comparative experiments on four benchmarks mentioned.
+As shown in Table 1,  10,  11 and  9, WePe achieves good detection performance on ImageNet, LSUN-BEDROOM, DRCT-2M and GenImage.
+Experimental results show the effectiveness of uncertainty estimation in detecting AI-generated images.
+It is worth noting that without any training, and relying only on the nature of the pre-trained model itself, WePe shows the potential to differentiate between natural and generated images.
+When further trained to amplify the uncertainty disparity between natural and generated images, WePe achieves superior average performance compared to other methods and demonstrates robust generalization capabilities across diverse datasets.
+To further illustrate the effectiveness of our method, we count the difference in feature similarity between natural and generated images on the pre- and post-perturbation models.
+As shown in Figure 4, small perturbation of the model has less effect on the natural images than on generated images, resulting in higher feature similarity before and after the perturbation.
+The discrepancy effectively distinguishes the natural image from the generated image.
+Comparison under attacks.
+In real-world scenarios, malicious actors may attempt to modify generated images to evade detection.
+To assess the robustness of the model under such conditions, we simulate an attack by introducing Gaussian noise (with a variance of $0.1$) to the generated images.
+Then the clean natural image and the attacked generated image are fed into the detector to determine whether the two can be reliably distinguished.
+Beyond spatial domain attacks, we also investigate attacks in the frequency domain, given the frequency differences between natural and generated images.
+As in Table 4, several detectors, such as NPR and RIGID, are vulnerable to this simple form of attack.
+In contrast, WePe, which identifies differences in the distributions of natural and generated images, demonstrates resilience to attacks, as the added noise further accentuates these distributional disparities, thereby enhancing WePe’s capacity to differentiate between the two types of images.
+Table 3: The effect of models.
+model AUROC    AP DINOv2: ViT-S/14   72.83   71.63 DINOv2: ViT-B/14   81.82   80.64 DINOv2: ViT-L/14   87.99   85.04 DINOv2: ViT-g/14   84.92   81.83 CLIP: ViT-L/14   84.82   84.20
 4.3 Ablation study
-In this section, we perform ablation experiments. Unless otherwise stated, experiments are conducted on ImageNet benchmark.
-Robustness to Perturbations.
-Robustness to various perturbations is a critical metric for detecting generated images.
+In this section, we perform ablation experiments. Unless otherwise stated, experiments are conducted on ImageNet.
+Robustness to Image Perturbations Robustness to various perturbations is a critical metric for detecting generated images.
 In real-world scenarios, images frequently undergo perturbations that can impact detection performance.
@@
 As illustrated in Figure 5, training-free methods generally exhibit superior robustness compared to training-based methods, with our method achieving the best overall performance.
+Selecting which layers’ parameters to perturb?
+As shown in Figure 7, we explore the choice of which layers’ parameters to perturb would achieve good performance.
+The horizontal coordinates in the graph indicate that the first $k$ blocks are perturbed, not the kthkth block.
+The experimental results exhibit that our method obtains good performance when the parameters of the first 9 to the first 20 blocks are chosen to be perturbed.
+This demonstrates the robustness of our method.
+In practice, we can select the layers to be perturbed by a small set of natural and generated images.
+And when the generated images are not available, we can also use the probe to determine which layers are perturbed using only the natural image.
+We describe our method in Appendix O.
+The effect of perturbation type.
+In experiments, model parameters are perturbed with Gaussian noise.
+We further explore other perturbation, such as adding uniform or Laplace noise to the weight.
+Besides, we also explore MC Dropout, i.e., using dropout during inference.
+As shown in Table 2, all three weight perturbation methods achieve good performance, and outperform MC Dropout.
 The impact of the degree of perturbation.
@@
 It is only when the noise is very large or very small that it leads to a degradation in performance.
-When the noise level is small, the features obtained before and after the model perturbation are extremely similar, while when the noise level is very large, the features obtained before and after the model perturbation are extremely dissimilar, and these two cases will result in the inability to effectively differentiate between real and generated images.
-Selecting which layers’ parameters to perturb?
-As shown in Figure 5, we explore the choice of which layers’ parameters to perturb would achieve good performance.
-The horizontal coordinates in the graph indicate that the first $k$ blocks are perturbed, not the $kth$ block.
-The experimental results exhibit that our method obtains good performance when the parameters of the first 9 to the first 20 blocks are chosen to be perturbed.
-This demonstrates the robustness of our method.
-In practice, we can select the layers to be perturbed by a small set of real and generated images.
-And when the generated images are not available, we can also use the probe to determine which layers are perturbed using only the real image.
-We describe our method in Appendix A.14.
+When the noise level is small, the features obtained before and after the model perturbation are extremely similar, while when the noise level is very large, the features obtained before and after the model perturbation are extremely dissimilar, and these two cases will result in the inability to effectively differentiate between natural and generated images.
 The effect of models.
@@
 We hypothesize that the difference comes from the training approach of these models.
-CLIP learns features using image captions as supervision, which may make the features more focused on semantic information, whereas DINOv2 learns features only from images, which makes it more focused on the images themselves, and thus better able to capture subtle differences in real and generated images.
-When using DINOv2, our methods performs poorly when the capacity of the model is small.
-This may be due to the fact that larger models can better capture the differences between real and fake images.
-In our main experiments, to balance detection efficiency and detection performance, we use DINOv2 ViT-L/14.
-The effect of perturbation type.
-In our experiments, we perturb the model parameters by adding Gaussian noise.
-We further explore other ways of perturbation, such as adding uniform noise or Laplace noise to the weight.
-In addition to this, we also explore the effect of MC Dropout, i.e., using dropout during inference.
-As shown in Table 5, all three weight perturbation methods achieve good performance, and outperform MC Dropout.
-Table 4: The effect of models.
-model      AUROC      AP DINOv2: ViT-S/14      72.83      71.63 DINOv2: ViT-B/14      81.82      80.64 DINOv2: ViT-L/14      87.99      85.04 DINOv2: ViT-g/14      84.92      81.83 CLIP: ViT-L/14      77.89      77.90
-Table 5: The effect of type of perturbation.
-noise      AUROC      AP Gaussian noise      87.99      85.04 Uniform noise      89.06      86.32 Laplace noise      87.13      84.22 MC Dropout      81.63      79.71
-5 Limitation
-The proposed weight perturbation provides a simple and effective method for detecting generated images, yet we have not theoretically justified the widespread use of the method due to the inclusion of a variety of strong prior assumptions, such as the assumption about treating generated samples as OOD data.
-Therefore, our future work will focus on establishing the theoretical foundations of our method.
+CLIP learns features using image captions as supervision, which may make the features more focused on semantic information, whereas DINOv2 learns features only from images, which makes it more focused on the images themselves, and thus better able to capture subtle differences in natural and generated images.
+5 Related work
+AI-Generated images detection.
+Recent advancements in generative models, such as those by (Brock et al.,, 2019; Ho et al.,, 2020), have led to the creation of highly realistic images, highlighting the urgent need for effective algorithms to distinguish between natural and generated images.
+Prior research, including works by (Frank et al.,, 2020; Marra et al.,, 2018), primarily focuses on developing specialized binary classification neural networks to differentiate between natural and generated images.
+Notably, CNNspot (Wang et al.,, 2020) demonstrates that a standard image classifier trained on ProGAN can generalize across various architectures when combined with specific data augmentation techniques.
+NPR (Tan et al.,, 2024) introduces the concept of neighboring pixel relationships to capture differences between natural and generated images.
+PatchCraft (Zhong et al.,, 2023) proposes an efficient AI-generated image detector by exploring texture patch artifacts.
+FatFormer (Liu et al.,, 2024) introduces a forgery-aware adaptive Transformer that adapts a pre-trained CLIP model to effectively discern and integrate local forgery traces from both image and frequency domains.
+DRCT (Chen et al.,, 2024) presents a Diffusion Reconstruction Contrastive Training framework to improve the generalizability of synthetic image detectors by training them to distinguish real images from their high-quality diffusion-based reconstructions.
+Although these methods show superior performance on generators in the training set, they often do not generalize well to unknown generators.
+In addition to this, training-based methods are susceptible to small perturbations in the image.
+For this reason, recently, some training-free methods have been proposed.
+AEROBLADE (Ricker et al.,, 2024) calculates the reconstruction error with the help of the autoencoder used in latent diffusion models (Rombach et al.,, 2022).
+RIGID (He et al.,, 2024) finds that natural images are more robust to small noise perturbations than generated images in the representation space of the vision foundation models and exploits this property for detection.
+However, these methods usually make overly strong assumptions about natural or generated images, leading to insufficient generalization.
+In our paper, we propose a training-free detection method through uncertainty analysis.
+Based on the widespread phenomenon that generated images have greater uncertainty than natural images on models trained with natural images, our method achieves robust detection performance.
+Uncertainty estimation.
+Uncertainty estimation in machine learning has seen significant advancements in recent years.
+(Gal and Ghahramani,, 2016) introduces Monte Carlo Dropout (MC Dropout), which uses dropout at inference to estimate uncertainty from the variance of multiple predictions.
+(Lakshminarayanan et al.,, 2017) develops deep ensembles, demonstrating improved uncertainty estimates through training multiple model independently with different initializations.
+Recent work by (Snoek et al.,, 2019) analyzes the calibration of uncertainty in deep learning models, highlighting the importance of reliable uncertainty measures.
+Additionally, (Guo et al.,, 2017) explore the use of temperature scaling to enhance the calibration of model predictions.
+(Blundell et al.,, 2015) introduce Bayes by Backprop, a method for estimating weight uncertainty in neural networks by modeling the posterior distribution over weights using variational inference, improving model generalization and robustness.
+Similarly, (Ferrante et al.,, 2024) leverage weight perturbation techniques to estimate neural network uncertainty, demonstrating improved classification accuracy through robust uncertainty quantification.
+(Pearce et al.,, 2020) explore distribution-free uncertainty estimation, using conformal prediction and quantile regression to estimate bounds on aleatoric uncertainty.
+(Chan et al.,, 2024) introduce hyper-diffusion models, allowing to accurately estimate both epistemic and aleatoric uncertainty with a single model.
 6 Conclusion
@@
 Future work could delve deeper into refining the predictive models and exploring additional features that could further enhance detection accuracy.
-ETHIC STATEMENT
-According to the Code of Ethics, our work does not raise any ethical concerns, as there are no human subjects, private datasets, harmful insights, or research integrity issues.
-Our work does not raise any ethical concerns.
-In particular, our work focuses on detecting AI-generated images to avoid the potential risk induced by the development of generative models.
-REPRODUCIBILITY STATEMENT
-To highlight the efforts that have been made to ensure reproducibility, we summarize materials facilitating reproducible results:
-  • Theoretical results. Our work mainly focuses on the empirical investigation without explicit theatrical results.
-  • Datasets. All involved datasets are publicly available. To facilitate reproducible results, we provide details in Sec 4 and Appendix A.15.
-  • Open Source. We will release our code once the paper is accepted. If reviewers request, we will include an anonymized link to the code in our response.
-Appendix A Appendix
-A.1 Concerns about models being contaminated by generated images
-With the proliferation of generated images, future large-scale models could become contaminated by such data, making it increasingly difficult to distinguish between natural and generated images.
-One potential solution is to employ machine unlearning (Yao et al., 2023) techniques to detect and address generated images.
-Machine unlearning focuses on removing the influence of specific data from pre-trained models, primarily due to privacy concerns.
-In this context, when generative images are incorporated into the training process of large-scale models, we can utilize machine unlearning techniques to eliminate the effects of these images on pre-trained models.
-This approach would help ensure that the features of natural and generated images remain distinct and separable.
-A.2 Discussion on distribution discrepancy
+Acknowledgments
+This work was supported in part by NSFC No.
+62222117.
+JN and BH were supported by NSFC General Program No.
+62376235, RGC Young Collaborative Research Grant No.
+C2005-24Y, RGC General Research Fund No.
+12200725, Guangdong Basic and Applied Basic Research Foundation Nos.
+2022A1515011652 and 2024A151501239, and HKBU CSD Departmental Incentive Scheme.
+TLL was partially supported by the following Australian Research Council projects: FT220100318, DP220102121,LP220100527,LP220200949.
+YGZ was funded by Inno HK Generative AI R&D Center.
+YCM was supported by the RGC Senior Research Fellow Scheme under the grant: SRFS2324-2S02.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes] ,
+  4. Justification: The main claims made in the abstract and introduction accurately reflects the paper’s contributions and scope.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes] ,
+  9. Justification: Limitations are discussed in Appendix A.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: The paper provides a full set of assumptions and a complete proof.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: The paper fully discloses all the information needed to reproduce the main experimental results.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The data and code will be released once prepared.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The paper provides details of implementation in Appendix Q.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We report the average results under five different random seeds and report the variance in Figure 10.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The paper provide sufficient information on the computer resources.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The paper has discussed societal impacts in Appendix B.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A] .
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have cited the original paper that produced the code package and dataset.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We communicate the details of the code as part of our submission via structured templates.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing and research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing and research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Limitation
+A key limitation of the proposed method lies in its reliance on the assumption of distinct uncertainty profiles between natural and generated images.
+As generative models continue to advance, the uncertainty gap between these image types, as captured by the foundational model, may progressively narrow.
+This convergence could undermine the method’s effectiveness, necessitating additional calibration techniques such as the proposed WePe, to amplify the differentiation in uncertainty.
+Appendix B Social impacts
+The proposed AI-generated image detection method contributes to mitigating societal risks posed by generative model fraud.
+By improving the ability to identify synthetic media, such as deepfakes, this work helps combat disinformation and enhances trust in digital content, particularly in critical domains like journalism and legal evidence.
+Appendix C Proof of THEOREM 3.2
+Proof.
+Consider a neural network $f(x;\theta):\mathcal{X}\to\mathbb{R}^{d}$, parameterized by $\theta\in\mathbb{R}^{p}$, which maps an input image $x$ to a feature vector $f(x;\theta)$.
+The model is trained on the natural image distribution $\mathcal{D}^{1}$ by minimizing a loss function:
+$\mathcal{L}(\theta)=\mathbb{E}_{x\sim\mathcal{D}^{1}}\left[\ell(f(x;\theta),y)\right],$ (17)
+where $\ell$ is the loss, and $y$ is the label. The optimal parameters are denoted $\theta^{*}$.
+To capture sensitivity, we define a loss function that measures the change in the feature vector under parameter perturbations:
+$\ell(x,\theta,\xi)=\|f(x;\theta+\xi)-f(x;\theta)\|_{2}^{2},$ (18)
+where $\xi\sim\mathcal{N}(0,\sigma^{2}I)$ represents a Gaussian perturbation with variance $\sigma^{2}$. For small $\xi$, we approximate:
+$f(x;\theta+\xi)\approx f(x;\theta)+\nabla_{\theta}f(x;\theta)^{\top}\xi,$ (19)
+so:
+$\ell(x,\theta,\xi)\approx\|\nabla_{\theta}f(x;\theta)^{\top}\xi\|_{2}^{2}=\xi^{\top}J(x)^{\top}J(x)\xi,$ (20)
+where $J(x)=\nabla_{\theta}f(x;\theta)$ is the Jacobian matrix. The expected loss over perturbations is:
+$\mathbb{E}_{\xi}[\ell(x,\theta,\xi)]\approx\mathbb{E}_{\xi}\left[\xi^{\top}J(x)^{\top}J(x)\xi\right]=\sigma^{2}\text{tr}\left(J(x)^{\top}J(x)\right)=\sigma^{2}\|J(x)\|_{F}^{2}.$ (21)
+Thus, the expected sensitivity is proportional to the expected loss:
+$\mathbb{E}_{x\sim\mathcal{D}}\left[\text{sen}(x)\right]=\mathbb{E}_{x\sim\mathcal{D}}\left[\|J(x)\|_{F}^{2}\right]=\frac{1}{\sigma^{2}}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{\xi}[\ell(x,\theta,\xi)]\geq 0.$ (22)
+According to PAC-Bayes theory (McAllester,, 1998), we consider a prior distribution $P$ over parameters $\theta$, typically $P=\mathcal{N}(\theta_{0},\sigma_{0}^{2}I)$, and a posterior distribution $Q$, typically $Q=\mathcal{N}(\theta^{*},\sigma^{2}I)$, where $\theta^{*}$ is the trained parameter.
+The PAC-Bayes theorem provides a bound on the expected loss under $Q$.
+For a loss function $\ell$, the standard PAC-Bayes bound states that, with probability at least $1-\delta$ over the draw of a training set $T=\{(x_{i},y_{i})\}_{i=1}^{N}\sim\mathcal{D}^{N}$, for any posterior $Q$:
+$\mathbb{E}_{\theta\sim Q}\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(\theta,x,y)]\leq\mathbb{E}_{\theta\sim Q}\left[\frac{1}{N}\sum_{i=1}^{N}\ell(\theta,x_{i},y_{i})\right]+\sqrt{\frac{\text{KL}(Q\|P)+\ln\frac{N}{\delta}}{2(N-1)}},$ (23)
+where $\text{KL}(Q\|P)$ is the Kullback-Leibler divergence between $Q$ and $P$.
+Here, we adapt the loss to $\ell(x,\theta,\xi)=\|f(x;\theta+\xi)-f(x;\theta)\|_{2}^{2}$. Since $\xi$ is a perturbation, we consider the expected loss over $\xi$, which can be formed as:
+$\ell^{\prime}(x,\theta)=\mathbb{E}_{\xi\sim\mathcal{N}(0,\sigma^{2}I)}\left[\|f(x;\theta+\xi)-f(x;\theta)\|_{2}^{2}\right]\approx\sigma^{2}\|J(x)\|_{F}^{2}.$ (24)
+Since $\ell^{\prime}(x,\theta)$ depends on $\theta$, we approximate by evaluating at $\theta^{*}$, and consider the empirical sensitivity on the training set $T\sim\mathcal{D}^{1}$:
+$\hat{\ell}^{\prime}(T,\theta)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\xi}\left[\|f(x_{i};\theta+\xi)-f(x_{i};\theta)\|_{2}^{2}\right].$ (25)
+According to Eq. (23), we have the following inequality:
+$\mathbb{E}_{\theta\sim Q}\mathbb{E}_{x\sim\mathcal{D}}[\ell^{\prime}(x,\theta)]\leq\mathbb{E}_{\theta\sim Q}[\hat{\ell}^{\prime}(T,\theta)]+\sqrt{\frac{\text{KL}(Q\|P)+\ln\frac{N}{\delta}}{2(N-1)}}.$ (26)
+For $\mathcal{D}=\mathcal{D}^{1}$, the training set $T\sim\mathcal{D}^{1}$, and the model is optimized at $\theta^{*}$. The empirical sensitivity is:
+$\hat{\ell}^{\prime}(T,\theta^{*})\approx\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\|J(x_{i})\|_{F}^{2}.$
+Since the model is optimized on $T\sim\mathcal{D}^{1}$, the training process minimizes the loss landscape’s curvature around $\theta^{*}$, leading to a flat minimum (Hochreiter and Schmidhuber,, 1997), which means the empirical sensitivity tends to 0.
+At the same time, as $n$ tends to infinity, resulting in the expectation sensitivity converging to the empirical sensitivity.
+Therefore, we have:
+$\mathbb{E}_{x\sim\mathcal{D}^{1}}\left[\sigma^{2}\|J(x)\|_{F}^{2}\right]\approx 0,$ (27)
+which is consistent with our experimental results observed in Figure 3.
+For $\mathcal{D}=\mathcal{D}^{0}$, we evaluate the expected sensitivity on generated images, but the training set remains $T\sim\mathcal{D}^{1}$.
+Since $\mathcal{D}^{0}$ is not optimized, the feature representations for $x\sim\mathcal{D}^{0}$ lie in regions of higher curvature in the loss landscape, leading to larger singular values of $J(x)$.
+According to Eq.
+(22) and Eq.
+(27), we have:
+$\mathbb{E}_{x\sim\mathcal{D}^{0}}\left[\sigma^{2}\|J(x)\|_{F}^{2}\right]|\geq 0\approx\mathbb{E}_{x\sim\mathcal{D}^{1}}\left[\sigma^{2}\|J(x)\|_{F}^{2}\right],$ (28)
+which implies $\mathbb{E}_{x\sim\mathcal{D}^{1}}\left[\text{sen}(x)\right]\leq\mathbb{E}_{x\sim\mathcal{D}^{0}}\left[\text{sen}(x)\right]$. Thus we complete this proof. ∎
+Appendix D Analysis of the Upper Bounds for Uncertainty Estimation
+In Eq.
+13, we employ an upper bound to estimate uncertainty.
+Next, we provide an analysis to establish the validity of this estimation.
+The inequality is obtained through the Cauchy-Schwarz inequality, and we can explore the validity of the upper bound by calculating the difference between the two sides of the inequality.
+We define the difference:
+[…]
+Thus, the difference $\triangle$ can be written as:
+$\triangle=\left\|f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j=1}^{n}f\left(\mathbf{x};\theta_{j}\right)\right\|^{2}\cdot\left\|f\left(\mathbf{x};\theta_{t}\right)\right\|^{2}-\left(\left\|f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j=1}^{n}f\left(\mathbf{x};\theta_{j}\right)\right\|\left\|f\left(\mathbf{x};\theta_{t}\right)\right\|\cos\theta\right)^{2}$ (30)
+where $\theta$ is the angle between $f\left(\mathbf{x};\theta_{t}\right)$ and $f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j}^{n}f\left(\mathbf{x};\theta_{j}\right)$.
+And finally, we obtain the following expression:
+$\triangle$ $=\left\|f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j=1}^{n}f\left(\mathbf{x};\theta_{j}\right)\right\|^{2}\cdot\left\|f\left(\mathbf{x};\theta_{t}\right)\right\|^{2}\left(1-\cos^{2}\theta\right)$ (31)
+$=\left\|f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j=1}^{n}f\left(\mathbf{x};\theta_{j}\right)\right\|^{2}\cdot\left\|f\left(\mathbf{x};\theta_{t}\right)\right\|^{2}\sin^{2}\theta$ (32)
+This result indicates that the smaller the angle between $f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j}^{n}f\left(\mathbf{x};\theta_{j}\right)$ and $f\left(\mathbf{x};\theta_{t}\right)$, the tighter the upper bound.
+In our experiment, the perturbation applied to the model was extremely small, causing the image features to remain virtually unchanged, resulting in an extremely weak deviation between $f\left(\mathbf{x};\theta_{k}\right)-\frac{1}{n}\sum_{j}^{n}f\left(\mathbf{x};\theta_{j}\right)$ and $f\left(\mathbf{x};\theta_{t}\right)$, thereby enabling the upper bound to provide a reliable basis for ranking the instances.
+Appendix E Performance under attacks.
+As presented in Table 4, we evaluate the robustness of the proposed method against malicious manipulations of generated images.
+Specifically, for the FDA and SDA scenarios, we apply zero-mean Gaussian noise with a standard deviation of 0.1 to perturb the generated images.
+The results demonstrate that our method exhibits strong robustness to these perturbations.
+Table 4: Performance on ImageNet under frequency domain attacks (FDA) and spatial domain attacks (SDA). We report AUROC.
+Models
+ADM ADMG LDM DiT BigGAN GigaGAN StyleGAN XL RQ-Transformer Mask GIT Average $\Delta$
+Methods FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA FDA SDA
+Training Methods
+CNNspot 59.47 52.15 59.88 51.86 58.41 51.05 55.43 51.29 71.85 53.87 58.54 50.36 60.27 51.52 54.74 48.71 59.45 50.32 59.78 51.24 -7.26 -15.80
+Ojha 85.16 85.28 84.09 83.67 85.94 84.97 85.14 84.33 92.17 94.77 85.39 90.18 88.16 86.73 85.22 86.38 88.71 91.75 86.17 87.56 +0.82 +2.21
+DIRE 53.82 52.93 51.19 54.89 53.18 50.33 54.17 49.37 52.59 50.02 55.54 51.95 52.44 50.44 54.62 48.15 52.11 51.29 53.30 51.04 +0.60 -1.66
+NPR 83.01 46.99 82.03 46.59 86.64 46.22 81.22 45.22 84.41 46.31 81.53 46.51 82.86 45.78 82.49 47.65 86.18 45.42 83.37 46.30 -2.63 -39.70
+WePe 96.42 93.86 93.67 88.58 94.73 84.90 91.47 81.93 99.64 96.67 98.54 94.69 99.46 97.33 98.49 95.64 98.67 91.79 96.79 91.71 +1.22 -3.86
+Training-free Methods
+AEROBLADA 40.30 23.08 42.87 27.87 45.51 28.18 43.93 28.77 41.84 27.09 42.70 29.05 47.01 32.78 51.62 31.96 51.62 32.71 44.90 29.05 -12.97 -28.82
+RIGID 80.79 35.45 73.01 34.37 65.66 33.83 63.09 33.37 77.99 35.60 73.91 35.32 73.23 34.58 83.01 36.19 76.62 34.36 74.15 34.77 -9.43 -48.81
+WePe 91.22 93.78 85.70 88.73 80.87 83.78 79.02 80.19 93.32 93.31 91.69 92.46 93.96 94.42 93.95 95.11 88.44 88.84 88.68 90.07 +0.69 +2.08
+Appendix F Discussion on distribution discrepancy
 In this paper, the core assumption we make is that there is data distribution discrepancy between natural and generated images.
-This assumption is valid for current generative models and has been confirmed by many works (Corvi et al., 2023; Tan et al., 2024; Ricker et al., 2024).
+This assumption is valid for current generative models and has been confirmed by many works (Tan et al.,, 2024; Ricker et al.,, 2024).
 This assumption is also the foundation of many generative image detection methods (we cannot distinguish between images that are indistinguishable).
@@
 However, this remains an observation, and we have not found theoretical proof despite reviewing the literature.
-We only observe a similar phenomenon in UnivFD (Ojha et al., 2023), where the feature distribution discrepancy is observed in the representation space of CLIP:ViT-L/14.
+We only observe a similar phenomenon in UnivFD (Ojha et al.,, 2023), where the feature distribution discrepancy is observed in the representation space of CLIP:ViT-L/14.
 That said, we can confirm the existence of feature distribution discrepancy of generated and natural images based on an important metric for evaluating generative models, the FID score.
-The FID score measures the feature distribution discrepancy between natural and generated images on the Inception network (Szegedy et al., 2015).
+The FID score measures the feature distribution discrepancy between natural and generated images on the Inception network (Szegedy et al.,, 2015).
 When the FID score is 0, it indicates that the two distributions do not differ.
@@
 $\bullet$ and $\bullet$ represent the feature of natural images and AI-generated images on the corresponding models.
-A.3 Measuring feature distribution discrepancy with FID scores
-We further use the ”FID” score to measure the difference in feature distribution between natural and generated images.
+Appendix G Measuring feature distribution discrepancy with FID scores
+We further use the "FID" score to measure the difference in feature distribution between natural and generated images.
 To avoid the effects of categories, we compute the FID scores using the DINOv2 model on the LSUN-BEDROOM benchmark.
@@
 AUROC 50.00 73.85 83.50 86.23 88.84 94.16 94.18 95.34
-A.4 Comparison with Gradient Cuff
-Gradient Cuff (Hu et al., 2024) focuses on detecting jailbreak attacks in Large Language Models (LLMs).
-It finds that the landscape of the refusal loss is more precipitous for malicious queries than for benign queries.
-And then it uses stochastic gradient estimation to estimate gradient and use the gradient norm as the decision score.
-Thus, we can also leverage this interesting work to identify the distribution discrepancy for detection.
-We use the similar way to estimate gradient as the decision score.
-As shown in Table 8, we surprisingly find that this method even surpasses WePe.
-This suggests that it is also possible to distinguish between natural and generated images by estimating the gradient.
-A.5 WePe on large multi-modal models
-In addition to CLIP, we further test the performance of WePe on BLIP (Li et al., 2022).
+Appendix H WePe on large multi-modal models
+In addition to CLIP, we further test the performance of WePe on BLIP (Li et al.,, 2022).
 As shown in Table 7, the performance of WePe is unsatisfactory on these multimodal models, which may be due to the fact that the image features of the multimodal models are more focused on semantic information, in line with our discussions.
-A.6 Performance on adversarial examples
-We further test WePe on adversarial examples.
-We simply add Gaussian noise (with different standard deviation $\sigma$) to the test samples to simulate the adversarial examples.
-We test three cases: adding noise on the natural image, adding noise on the generated image and adding noise on all images.
-As shown in Table 10, when noise is injected, the feature similarity between the clean model and the noisy model for the noisy image decreases, which leads to a change in the detection performance.
-To mitigate this effect, we can perform detection by perturbing the model multiple times and using the average similarity.
-The effect of noise is successfully mitigated by ensemble as shown in Table 10.
-Table 7: Comparison with Gradient Cuff.
-Method      AUROC      AP WePe      87.99      85.04 WePe + Gradient Cuff      89.36      90.62
-Table 8: WePe on large multi-modal models.
-Model      AUROC      AP DINOv2      87.99      85.04 CLIP      77.89      77.90 BLIP      68.25      64.68
-Table 9: Performance on adversarial examples.
-Case      AUROC      AP Clean images      87.99      85.04 Noisy natural images, $\sigma=0.05$      81.79      78.99 Noisy natural images, $\sigma=0.1$      76.16      73.16 Noisy natural images, $\sigma=0.15$      69.53      66.83 Noisy generated images, $\sigma=0.05$      88.03      85.82 Noisy generated images, $\sigma=0.1$      88.77      87.14 Noisy generated images, $\sigma=0.15$      90.25      88.99 Noisy natural and generated images, $\sigma=0.05$      84.94      81.98 Noisy natural and generated images, $\sigma=0.1$      81.46      78.25 Noisy natural and generated images, $\sigma=0.15$      78.19      75.32
-Table 10: Mitigating the effects of image noise through ensemble.
-Case      AUROC      AP Clean images      87.99      85.04 Noisy natural images, $\sigma=0.05,n=1$      81.79      78.99 Noisy natural images, $\sigma=0.05,n=5$      85.30      82.59 Noisy natural images, $\sigma=0.1,n=1$      76.16      73.16 Noisy natural images, $\sigma=0.1,n=5$      81.36      78.39 Noisy natural images, $\sigma=0.15,n=1$      69.53      66.83 Noisy natural images, $\sigma=0.15,n=5$      75.16      72.58
-A.7 Concerns about hard samples
-WePe relies on the model being pre-trained on a large dataset of natural images.
-Given the abundance of natural images, it is possible WePe may misclassify other natural images that are out-of-distribution, as AI-generated images, leading to false-negative errors.
-Thus, we follow previous work to leverage AUROC as the main metric the evaluate different methods.
-This is because AUROC is a pivotal metric for reflecting the false-positive and false-negative costs.
-The hard (natural) samples play a crucial role in detecting generated images.
-We will explore how to leverage hard samples to promote the detection performance in our future work.
-A.8 Boosting performance on CLIP
-As mentioned above, WePe does not perform satisfactorily on vision language models such as CLIP.
-Since CLIP needs to unite image features and text features, the image features need to be projected into the union space, which results in the projected image features being more focused on semantic information.
-For this reason, we can improve the performance of WePe on CLIP by using the features before projection for detection, as shown in Table 12.
-Table 11: Boosting performance on CLIP.
-Model      AUROC      AP DINOv2      87.99      85.04 CLIP      77.89      77.90 CLIP without projection matrix      84.82      84.20
 Table 6: Comparison of detection times.
 Method      Time (s) AEROBLADE      17.6 RIGID      3.7 WePe      4.5
-A.9 Comparison of detection times.
-Table 13: Evading detection of RIGID by adding noise to the generated image.
-Model      AUROC      AP
-RIGID      83.58      81.58
-WePe      87.99      85.04
-RIGID with noisy generated images      18.69      34.51
-WePe with noisy generated images      88.77      87.14
+Table 7: WePe on large multi-modal models.
+Model      AUROC      AP DINOv2      87.99      85.04 BLIP      68.25      64.68
+Appendix I Comparison of computational costs.
 Our method use a perturbed pre-trained model that is fixed during inferring all test samples.
@@
 However, WePe can be parallelized across two devices to obtain the detection results in a single forward pass.
-A.10 Comparison with RIGID
-The main differences between WePe and RIGID are as follows:
-  • The approach proposed by RIGID stems only from the phenomenon it observed: namely, that natural and generated images show different sensitivities to noise in the representation space of DINOv2.
-Instead, WePe explicitly proposes that there is distribution discrepancy between natural and generated images, and utilizes the difference in uncertainty to expose distribution discrepancy between natural and generated images.
-  • RIGID utilizes the difference in sensitivity to noise between natural image and generated image for detection.
-Although the generated image is more sensitive, it is easy to think of a way to avoid detection, i.e., the user adds noise to the generated image and then submits it for detection.
-This approach easily removes the sensitivity of the generated image to noise.
-As shown in Table 13, after adding noise to a generated image, RIGID determines the noisy generated image as natural images.
-This is a fatal flaw of RIGID.
-However, WePe is not affected by this.
-WePe exposes the difference in distribution between the test image and the natural image through weight perturbation.
-Adding noise to the generated image further increases this difference, leading to even better detection performance.
-A.11 Note on the unbiased assumption
-In section 3.3, we make an assumption that the expected extracted feature by noised models is unbiased for that by the original model.
-We think this assumption is reasonable.
-Thanks to the over-parameterization of modern neural networks and advanced optimization algorithms (e.g., AdamW), it is a well-established fact that trained neural networks are usually smooth in the parameter space and show robustness to small weight perturbation (Novak et al., 2018).
-And this robustness is used in many applications, such as quantization (Gholami et al., 2022) and pruning (Liu et al., 2019).
-And in Figure 2, we also clearly show this robustness: when adding tiny noise to the model weights, the features remain almost unchanged.
-A.12 SOFTWARE AND HARDWARE
-We use python 3.8.16 and Pytorch 1.12.1, and seveal NVIDIA GeForce RTX-3090 GPU and NVIDIA GeForce RTX-4090 GPU.
-Figure 8: WePe with multiple perturbations.
-A.13 WePe with multiple perturbation
+In practice, WePe has high computational efficiency and is suitable for large-scale applications.
+Appendix J WePe with multiple perturbation
 In our experiments, taking into account the detection efficiency, we perturb the model only once, and then calculate the feature similarity of the test samples on the clean and perturbed models.
 We further experiment with multiple perturbations and use the mean of the feature similarity of the test samples on the clean model and all the perturbed models as the criterion for determining whether or not the image is generated by the generative models.
-As shown in Figure 8,
-A.14 Using natural images only to select which layers to perturb
+As shown in Figure 10, multiple perturbations can further improve performance.
+Table 8: The results of using classification models.
+model score AUROC    AP ResNet18 MSP   48.85   49.23 ResNet18 entropy   51.76   50.28 ViT-L/16 MSP   63.23   60.16 ViT-L/16 entropy   65.79   61.97
+Appendix K Comparison with OOD detection and uncertainty quantification methods
+  • Contrast with OOD Detection: Conventional OOD detection (Sun and Li,, 2022; Djurisic et al.,, 2023; Nie et al.,, 2024) usually relies on Maximum Softmax Probability (MSP) scores, exploiting fixed ID categories to identify OOD samples with low probabilities.
+In contrast, AI-generated detection involves diverse, unbounded categories, rendering MSP scores ineffective.
+WePe introduces a novel uncertainty estimation approach tailored for detecting generated images, achieving robust performance where traditional OOD methods falter.
+As shown in Table 8, we further use the ImageNet pre-trained classification model and used MSP and entropy as the scoring function to evaluate their performance on the AI-generated image detection task.
+The results show that these methods fail.
+  • Contrast with Uncertainty Quantification: Standard techniques like MC-Dropout and DeepEnsemble are ill-suited for DINOv2.
+Training multiple DINOv2 models for DeepEnsemble is computationally infeasible, and the absence of dropout in DINOv2 undermines MC-Dropout’s efficacy.
+Our proposed weight perturbation method overcomes these limitations, delivering a practical and effective uncertainty estimation tailored to DINOv2’s architecture, as validated by our experiments.
+Appendix L Analysis of Failure Cases
+WePe leverages the differential epistemic uncertainty exhibited by pre-trained models, such as DINOv2, when processing natural versus generated images to differentiate between them.
+Specifically, DINOv2, having been trained on an extensive dataset of natural images, demonstrates lower epistemic uncertainty for such images.
+Images that diverge from the distribution of the training dataset—despite not being generated by a generative model—tend to elicit higher uncertainty from DINOv2, leading to their misclassification as generated images.
+As shown in Figure 9, we visualized the feature shifts in DINOv2’s representations for a set of cartoon images, which were not produced by generative models, before and after weight perturbation.
+These images, due to their deviation from the training natural distribution, exhibited significant feature shifts post-perturbation, rendering them indistinguishable from generated images in our analysis.
+Figure 9: Feature shifts after model perturbation.
+Images are sampled from the following distributions: (a) natural image distribution, (b) AI-generated image distribution, and (c) cartoon image distribution.
+Appendix M Experiment results on GenImage, LSUN-BEDROOM and DRCT-2M
+As shown in Table 9, Table 10 and Table 11, our method achieves good performance on GenImage LSUN-BEDROOM and DRCT-2M, confirming the robustness of the proposed method.
+Table 9: AI-generated image detection performance (ACC, %) on GenImage.
+Methods Midjourney SD V1.4 SD V1.5 ADM GLIDE Wukong VQDM BigGAN Average
+Training Methods
+ResNet-50 54.9 99.9 99.7 53.5 61.9 98.2 56.6 52.0 72.1
+DeiT-S 55.6 99.9 99.8 49.8 58.1 98.9 56.9 53.5 71.6
+Swin-T 62.1 99.9 99.8 49.8 67.6 99.1 62.3 57.6 74.8
+CNNspot 52.8 96.3 95.9 50.1 39.8 78.6 53.4 46.8 64.2
+Spec 52.0 99.4 99.2 49.7 49.8 94.8 55.6 49.8 68.8
+F3Net 50.1 99.9 99.9 49.9 50.0 99.9 49.9 49.9 68.7
+GramNet 54.2 99.2 99.1 50.3 54.6 98.9 50.8 51.7 69.9
+DIRE 60.2 99.9 99.8 50.9 55.0 99.2 50.1 50.2 70.7
+UnivFD 73.2 84.2 84.0 55.2 76.9 75.6 56.9 80.3 73.3
+PatchCraft 79.0 89.5 89.3 77.3 78.4 89.3 83.7 72.4 82.3
+NPR 81.0 98.2 97.9 76.9 89.8 96.9 84.1 84.2 88.6
+FatFormer 92.7 100.0 99.9 75.9 88.0 99.9 98.8 55.8 88.9
+GenDet 89.6 96.1 96.1 58.0 78.4 92.8 66.5 75.0 81.6
+DRCT 91.5 95.0 94.4 79.4 89.1 94.6 90.0 81.6 89.4
+WePe 91.7 99.5 98.4 82.3 93.6 98.1 95.0 87.1 93.2
+Training-free Methods
+AEROBLADE 80.3 87.5 86.8 67.2 81.5 83.7 51.1 52.5 73.83
+RIGID 81.54 69.5 68.72 72.35 84.15 68.57 78.98 93.02 78.19
+WePe 79.17 77.8 75.57 76.07 79.20 79.00 90.60 89.27 80.84
+Table 10: AI-generated image detection performance on LSUN-BEDROOM.
+Models
+ADM DDPM iDDPM Diffusion GAN Projected GAN StyleGAN Unleashing Transformer Average
+Methods AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP
+CNNspot 64.83 64.24 79.04 80.58 76.95 76.28 88.45 87.19 90.80 89.94 95.17 94.94 93.42 93.11 84.09 83.75
+Ojha 71.26 70.95 79.26 78.27 74.80 73.46 84.56 82.91 82.00 78.42 81.22 78.08 83.58 83.48 79.53 77.94
+DIRE 57.19 56.85 61.91 61.35 59.82 58.29 53.18 53.48 55.35 54.93 57.66 56.90 67.92 68.33 59.00 58.59
+NPR 75.43 72.60 91.42 90.89 89.49 88.25 76.17 74.19 75.07 74.59 68.82 63.53 84.39 83.67 80.11 78.25
+WePe 79.41 76.68 96.71 96.16 94.18 93.44 99.81 99.80 99.83 99.82 97.06 96.51 99.45 99.37 95.21 94.54
+AEROBLADA 57.05 58.37 61.57 61.49 59.82 61.06 47.12 48.25 45.98 46.15 45.63 47.06 59.71 57.34 53.85 54.25
+RIGID 71.90 72.29 88.31 88.55 84.02 84.80 91.42 91.90 92.12 92.54 77.29 74.96 91.37 91.39 85.20 85.20
+WePe 73.85 70.21 88.84 87.14 86.23 83.82 94.16 93.52 95.34 95.18 83.50 80.66 94.18 93.45 88.01 86.28
+Table 11: AI-generated image detection performance (ACC, %) on DRCT-2M.
+Method SD Variants Turbo Variants LCM Variants ControlNet Variants DR Variants Avg.
+LDM SDv1.4 SDv1.5 SDv2 SDXL
+SDXL-
+Refiner
+SD-
+Turbo
+SDXL-
+Turbo
+LCM-
+SDv1.5
+LCM-
+SDXL
+SDv1-
+Ctrl
+SDv2-
+Ctrl
+SDXL-
+Ctrl
+SDv1-
+DR
+SDv2-
+DR
+SDXL-
+DR
+CNNSpot 99.87 99.91 99.90 97.63 66.25 86.55 86.15 72.42 98.26 61.72 97.96 85.89 82.94 60.93 51.41 50.28 81.12
+F3Net 99.85 99.78 99.79 88.60 55.85 87.37 63.29 63.66 97.39 54.98 97.98 72.39 81.99 65.42 50.39 50.27 71.13
+CLIP/RN50 99.00 99.99 99.96 94.61 62.08 91.43 84.40 64.40 98.97 57.43 99.74 80.69 82.03 65.83 50.67 50.47 80.05
+GramNet 99.40 99.01 98.84 95.30 62.63 80.68 71.19 69.32 93.05 57.02 89.97 75.55 82.68 51.23 50.01 50.08 76.62
+De-fake 92.10 95.53 99.51 89.65 64.02 69.24 92.00 93.93 99.13 70.89 58.98 62.34 66.66 50.12 50.16 50.00 75.52
+Conv-B 99.97 100.0 99.97 95.84 64.44 82.00 60.75 99.27 99.27 62.33 99.80 83.40 73.28 61.65 51.79 50.41 79.11
+Ojha 98.30 96.22 96.33 93.83 91.01 93.91 86.38 85.92 90.44 89.99 90.41 81.06 89.06 51.96 51.03 50.46 83.46
+DIRE 54.62 75.89 76.04 99.87 59.90 93.08 97.55 87.29 72.53 67.85 99.69 64.40 64.40 49.96 52.48 49.92 72.55
+DRCT 94.45 94.35 94.24 95.05 96.41 95.38 94.81 94.48 91.66 95.54 93.86 93.50 93.54 84.34 83.20 67.61 91.35
+FatFormer 96.52 95.31 93.27 91.99 92.87 91.78 88.15 87.48 92.82 91.76 90.28 86.99 88.19 65.92 60.15 55.13 85.53
+WePe 92.38 67.18 65.88 74.05 75.62 72.23 66.82 62.46 66.88 77.25 75.41 74.92 80.34 63.98 59.65 59.68 70.92
+WePe 97.06 96.03 94.76 96.45 96.59 97.81 93.54 92.66 96.29 94.43 96.69 96.17 95.72 75.99 73.32 69.78 91.45
+Appendix N SOFTWARE AND HARDWARE
+We use python 3.8.16 and Pytorch 1.12.1, and several NVIDIA GeForce RTX-3090 GPU and NVIDIA GeForce RTX-4090 GPU.
+Figure 10: WePe with multiple perturbations.
+Appendix O Using natural images only to select which layers to perturb
 In our experiments, we use a small set of natural images and generated images to pick the parameters that need to be perturbed.
-When all the generated images are not available, we can also use only the real images to select the layers that need to be perturbed.
+When all the generated images are not available, we can also use only the natural images to select the layers that need to be perturbed.
 Specifically, we first perturb each block alone and calculate the similarity of the features on the model of the natural image before and after the perturbation, as shown in Table 12.
@@
 WePe top-18 90.02 87.83 83.39 80.58 79.12 74.64 76.18 71.12 91.82 91.36 92.26 91.71 93.77 93.39 93.68 92.89 89.12 87.57 87.71 85.68
-A.15 Details of Datasets
+Appendix P Details of Datasets
 IMAGENET.
-The real images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval.
-The images are provided by (Stein et al., 2023).
+The natural images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval.
+The images are provided by (Stein et al.,, 2023).
 The generative model includes: ADM, ADMG, BigGAN, DiT-XL-2, GigaGAN, LDM, StyleGAN-XL, RQ-Transformer and Mask-GIT.
-The resolution of real images and generated images are $256\times 256$.
-We crop the image randomly to $224\times 224$ resolution.
 LSUN-BEDROOM.
-The real images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval.
-The images are provided by (Stein et al., 2023).
+The natural images and generated images can be obtained at https://github.com/layer6ai-labs/dgm-eval.
+The images are provided by (Stein et al.,, 2023).
 The generative model includes: ADM, DDPM, iDDPM, StyleGAN, Diffusion-Projected GAN, Projected GAN and Unleashing Transformers.
-The resolution of real images and generated images are $256\times 256$.
-We crop the image randomly to $224\times 224$ resolution.
-GenImagThe real images and generated images can be obtained at https://github.com/GenImage-Dataset/GenImage.
-The images are provided by (Zhu et al., 2023).
+GenImage.
+The natural images and generated images can be obtained at https://github.com/GenImage-Dataset/GenImage.
+The images are provided by (Zhu et al.,, 2023).
 The generative model includes: Midjourney, SD V1.4, SD V1.5, ADM, GLIDE, Wukong, VQDM and BigGAN.
-The real images come from ImageNet, and different images have different resolutions.
-Following (Stein et al., 2023), we resize the image to $256\times 256$ resolution and adjust its format to keep the same with the generated images, then we randomly crop it to $224\times 224$ resolution to extract features.
-For the generated images, we report our processing in detail as follows:
-  • Midjourney. The resolution of images generated by Midjourney is $1024\times 1024$, and we randomly crop them to $224\times 224$ resolution.
-  • SD V1.4. The resolution of images generated by SD V1.4 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
-  • SD V1.5. The resolution of images generated by SD V1.5 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
-  • ADM. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
-  • GLIDE. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
-  • Wukong. The resolution of images generated by SD V1.5 is $512\times 512$, and we randomly crop them to $224\times 224$ resolution.
-  • VQDM. The resolution of images generated by SD V1.5 is $256\times 256$, and we randomly crop them to $224\times 224$ resolution.
-  • BigGAN. The resolution of images generated by SD V1.5 is $128\times 128$, and we fill them with zero pixels to $224\times 224$ resolution.
+The natural images come from ImageNet, and different images have different resolutions.
+DRCT-2M.
+The natural images of DRCT-2M come from CoCo and can be obtained from https://cocodataset.org/#download.
+AI-generated images of DRCT-2M can be obtained from https://modelscope.cn/datasets/BokingChen/DRCT-2M/files, which are provided by (Chen et al.,, 2024).
+The generative model includes LDM, SDv1.4, SDv1.5, SDv2, SDXL, SDXL-Refiner, SD-Turbo, SDXL-Turbo, LCM-SDv1.5, LCM-SDXL, SDv1-Ctrl, SDv2-Ctrl, SDXL-Ctrl, SDv1-DR, SDv2-DR, SDXL-DR.
+Appendix Q Implementation details
+To balance detection performance and efficiency, we use DINOv2 ViT-L/14.
+We report the average results under five different random seeds and report the variance in Figure 10.
+In our experiments we find that perturbing the high layers may lead to a large corruption in the features of the natural images, resulting in sub-optimal results.
+Therefore, We do not perturb the high-level parameters.
+In DINOv2 ViT-L/14, the model has 24 transformer blocks, and we only perturb the parameters of the first 19 blocks with Gaussian perturbations of zero mean.
+The variance of Gaussian noise is proportional to the mean value of the parameters in each block, with the ratio set to 0.1.
+Considering the computational cost, we perturb the model only $1$ time, i.e., $n=1$.
+Multiple perturbations can further improve the performance as shown in Table 10.
+For WePe, we leverage LoRa (Hu et al.,, 2022) for parameter-effcient fine-tuning.
+The Lora layers are applied on the q_proj and v_proj layers of DINOv2.
+lora_rlora\_r and $lora\_\alpha$ are set to 8.
+And the model is optimized using the AdamW optimizer with a learning rate of $1\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a weight decay of $0.01$.
+Following CNNspot (Wang et al.,, 2020), data augmentation techniques including JPEG compression and Gaussian blur are employed to enhance robustness.
+For the IMAGENNET and LSUN-BEDROOM benchmarks, the ProGAN dataset serves as the training set.
+For the GenImage benchmark, SDv1.4 dataset is used as training set.
+For the DRCT-2M benchmark, SDv2 dataset is used as training set.
+When testing, to ensure objectivity in calculating classification accuracy and mitigate biases arising from manually selected thresholds, following (Ojha et al.,, 2023), we automatically determine the optimal threshold by identifying the score that maximizes the separation between natural and generated images, based on their computed classification scores.